### v1

# A Graph-Theoretical Perspective on Law Design for Multiagent Systems

Abstract
A law in a multiagent system is a set of constraints imposed on agents’ behaviours to avoid undesirable outcomes. The paper considers two types of laws: useful laws that, if followed, completely eliminate the undesirable outcomes and gap-free laws that guarantee that at least one agent can be held responsible each time an undesirable outcome occurs. In both cases, we study the problem of finding a law that achieves the desired result by imposing the minimum restrictions.
We prove that, for both types of laws, the minimisation problem is NP-hard even in the simple case of one-shot concurrent interactions. We also show that the approximation algorithm for the vertex cover problem in hypergraphs could be used to efficiently approximate the minimum laws in both cases.
1 Introduction
Suppose that three factories $a$, $b$, and $c$ need to dump the same type of pollutant into a river. Each factory must dump once every three days. The assimilative capacity of the river endures at most two factories dumping per day. Otherwise, the fish in the river would be killed. To avoid the death of the fish, the local government would like to set a law that regulates the dumping activity. Building on the well-known legal maxim “everything which is not forbidden is allowed”, we assume that a law serves solely to identify the actions from which agents must abstain. This is in line with the liberal rule-of-law perspective: a key virtue of the rule of law is the protection of individual freedom (hayek1944road; raz1979rule).
In a simple form, a dumping law can assign each factory a fixed dumping day within a recurring three-day cycle by banning it from dumping on the other two days. For instance, the law specifies the set $L_{0}=\{d_{a}^{1},d_{a}^{2},d_{b}^{2},d_{b}^{3},d_{c}^{1},d_{c}^{3}\}$ of banned dumping actions, where $d_{x}^{i}$ represents the action that factory $x$ dumping on the $i^{\text{th}}$ day in each three-day cycle. Under this law, factory $a$ dumps on the third day ($d_{a}^{3}$), factory $b$ dumps on the first day ($d_{b}^{1}$), and factory $c$ dumps on the second day ($d_{c}^{2}$) of each three-day cycle. In other words, only one factory dumps on each day, and it is guaranteed that the fish will not be killed. In this case, we say that law $L_{0}$ is useful in terms of prohibiting the death of the fish. In general, we say that a law is useful if the prohibited outcomes shall never appear when every agent obeys the law. The term “useful” is adopted from (shoham1995social), a pioneering work on law design in multiagent systems.
It is easily observable that the law $L_{0}$ above unnecessarily constrains the dumping behaviour of the factories. Note that, to avoid the death of the fish, it suffices to ensure that not all three factories dump on the same day. In other words, for any given day, the law only needs to prevent one factory from dumping. This means that the set $L_{1}=\{d_{a}^{1},d_{b}^{2},d_{c}^{3}\}$ is also a useful law. However, law $L_{1}$ allows each factory one more day to dump in each three-day cycle, providing more flexibility toward the factories’ production activities.
Observe that laws $L_{0}$ and $L_{1}$ are both useful, while $L_{1}$ sets fewer constraints than $L_{0}$ (i.e. $L_{1}\subsetneq L_{0}$). In this case, we say that $L_{1}$ is a useful reduction of law $L_{0}$. Also, notice that any further reduction of law $L_{1}$ is no longer useful. For example, the reduced law $L_{2}=\{d_{a}^{1},d_{b}^{2}\}$ of law $L_{1}$ allows all factories to dump on the third day and kill the fish. In other words, law $L_{1}$ satisfies a minimality property regarding usefulness. In general, we say that a law is minimal-useful if it is useful but cannot be further reduced while keeping usefulness. Considering the minimality of law captures the idea of setting minimal constraints on society, which is in line with the opinion that “the minimal state is the most extensive state that can be justified. Any state more extensive violates people’s rights …” (nozick1974anarchy).
Essentially, the usefulness of a law captures the ability of the law to prevent the undesirable outcomes. This is a typical focus in the literature on law design in multiagent systems. While reliable, the usefulness requirement excludes the possibility of coordination among agents as a means of prevention. In our example, even without a law, the three factories can still negotiate a dumping plan to avoid the death of the fish. On one hand, a negotiated plan could be more adaptive to the production of the factories than a law and thus bring higher efficiency. On the other hand, a law should be stable (jefferson1787unstability; raz1979rule). In contrast, coordination achieved by negotiation is flexible and thus can better accommodate changes in production. Although appealing, coordination is not always realisable in multiagent settings. In our case, the three factories may not be able to reach an agreement on the dumping plan due to the lack of communication or the conflict of interest. In general, coordination is even harder to achieve in complex multiagent systems, particularly when different types of agents exist (e.g. a traffic system including autonomous vehicles and human drivers, or a ranch including sheep, sheepdogs, herders and wolves).
In this paper, we relax the usefulness requirement and consider laws that can simultaneously accommodate coordination and the failure of coordination. Informally, under such laws, prohibited outcomes may appear when the agents all obey the law but do not coordinate. Nevertheless, there is always at least one principal agent who has a safe action that is lawful, and enables the prevention of prohibited outcomes without the need for coordination. As an example, law $L_{2}$ satisfies the above property. Recall that law $L_{2}$ only bans factory $a$ from dumping on the first day and factory $b$ from dumping on the second day. Consequently, even if all factories comply with law $L_{2}$, without coordination, the death of the fish may still happen when they simultaneously dump on the third day. However, factory $c$ has a safe action under law $L_{2}$ (i.e. dumping on either the first or the second day) that can solely prevent the death of the fish as long as factories $a$ and $b$ obey law $L_{2}$. By this means, law $L_{2}$ leaves factory $c$ a chance to dump on the third day while keeping the fish alive through negotiation with other factories. Meanwhile, if the negotiation fails, factory $c$ can still prevent the death of the fish on its own. In other words, factory $c$ is a principal agent under law $L_{2}$.
Observe that, under such laws, if a prohibited outcome finally appears, then either there is an agent whose action breaks the law, or the principal agent fails to prevent it. In the former case, we say that the agent who breaks the law bears legal responsibility. In the latter case, we say that the principal agent who could have prevented the prohibited outcomes with a safe action but fails to do so bears counterfactual responsibility. In our example, if the fish is killed because all three factories dump on the first day, then factory $a$ breaks law $L_{2}$ and thus $a$ is legally responsible; if it happens on the second day, then factory $b$ breaks law $L_{2}$ and thus $b$ is legally responsible; if it happens on the third day, then factory $c$, the principal agent under law $L_{2}$, fails to utilise her safe action to prevent it, and thus $c$ is counterfactually responsible. In a word, under those laws, if a prohibited outcome happens, then there is at least one agent either legally or counterfactually responsible. That is, a responsible agent can always be identified for any prohibited outcome. In this sense, we refer to such laws responsibility gap-free (abbr. gap-free).
To further clarify our terminology, counterfactual responsibility is usually regarded as a form of moral responsibility (robb2023moral). It captures the principle of alternative possibilities (frankfurt1969alternate): a person is morally responsible for what she has done only if she could have done otherwise. In the recent literature on responsibility in multiagent systems, the component “could have done otherwise” is commonly interpreted as an agent’s strategic ability to guarantee prevention irrespective of the behaviour of other agents (naumov2019blameworthiness; yazdanpanah2019strategic; baier2021game; shi2024responsibility). Note that the former discussion about the principal agent’s safe action to prevent prohibited outcomes aligns with such an interpretation. In this sense, our use of the term counterfactual responsibility in this paper (i.e. principal agents who fail to prevent are counterfactually responsible) is consistent with its treatment in the existing literature. As for responsibility gap, also called responsibility void, it is one of the important topics discussed in the ethics literature, especially in the context of artificial intelligence (matthias2004responsibility; braham2011responsibility; duijf2018responsibility; braham2018voids; burton2020mind; gunkel2020mind; langer2021we; goetze2022mind). Informally, responsibility gap captures the situation where an undesired outcome occurs but no agent can be held responsible, which is usually regarded as “unwanted” (hiller2022axiomatic). Correspondingly, the term gap-freeness in this paper precisely captures the absence of responsibility gaps as discussed in the literature and serves as a desired property of multiagent systems.
Back to the point, by the former analysis, law $L_{2}$ is gap-free. It sets fewer constraints on the agents’ actions than the minimal-useful law $L_{1}$ and allows for potential coordination. However, law $L_{2}$ is not minimal yet in terms of gap-freeness. Let us consider the law $L_{3}=\{d_{a}^{1}\}$, a reduction of law $L_{2}$. Since law $L_{3}$ bans factory $a$ from dumping on the first day, as long as factory $b$ dumps on that day (i.e. safe action), the fish would survive regardless of factory $c$’s choice. The same applies to factory $c$. In other words, both factories $b$ and $c$ are principal agents under law $L_{3}$. In addition, under law $L_{3}$, if the fish is killed because all factories dump on the first day, then factory $a$ is legally responsible; if it happens on the second or the third day, then factories $b$ and $c$ are both counterfactually responsible. This implies the gap-freeness of law $L_{3}$. On top of it, we say that law $L_{3}$ is a gap-free reduction of law $L_{2}$ because $L_{3}\subsetneq L_{2}$. Moreover, observe that the law $L_{4}=\varnothing$ is not gap-free because the death of the fish may happen, but no agent has a safe action that can individually prevent it under law $L_{4}$. This shows that any further reduction of law $L_{3}$ is not gap-free, and thus law $L_{3}$ satisfies a minimality property in terms of gap-freeness. In general, we say that a law is minimal-gap-free if it is gap-free but cannot be further reduced while keeping gap-freeness.
Contribution
In this paper, we investigate the design of useful laws and gap-free laws in multiagent systems. In particular, we model multiagent systems as one-shot concurrent games and interpret the law design problem in a graph-theoretical perspective. Specifically, we first formalise the concepts of usefulness and gap-freeness in Section 2. Then, in Section 3, we establish the equivalence between the useful law design problem and the vertex cover problem in hypergraphs (a.k.a. hitting set problem) by providing polynomial-time reductions between them. After that, in Section 4, we show that the gap-free law design problem is at least as hard as the useful law design problem, and further present a method for solving the gap-free law design problem by reducing it to the vertex cover problem in hypergraphs. Note that the vertex cover problem is one of the most important problems in complexity theory, listed as one of the 21 NP-complete problems by karp1972reducibility. Extensive research has focused on its hardness and approximation (chvatal1979greedy; bar1985local; slavik1996tight; feige1998threshold).
Novelty
By reducing the law design problems to the vertex cover problem, we make it possible to tackle the computational intractability in the law design problems using approximation techniques. This is distinct from the literature on law design (i.e. norm synthesis) in multiagent systems. In fact, there has been a rich amount of studies about normative system, where laws/norms are used to regulate the multiagent systems. They mainly use first-order logic (shoham1995social; fitoussi2000choosing) or modal logic (van2007social; aagotnes2009temporal) to describe the concerned properties of a system and use deontic logic to capture laws/norms (alechina2018norm). In this way, they are capable of modelling more complex systems than one-shot concurrent games. Meanwhile, they study both offline design (i.e. design-time norm synthesis) (fitoussi2000choosing; aagotnes2012conservative) and online design (i.e. run-time norm synthesis) (morales2011using; morales2013automated; morales2014minimality; riad2022run), considering both static and dynamic norms that may evolve (alechina2022automatic; riad2022run). Depending on the difference in object and formalisation, the computational complexity of their problems ranged from NP-complete (shoham1995social) to beyond EXPTIME (perelli2019enforcing; galimullin2024synthesizing). Although few of them try to address the complexity by heuristics (christelis2009automated) or by degenerating their problems into optimisation problems (aagotnes2010optimal; wu2022bayesian), none have tackled the intractability by considering inapproximability in addition to approximation, as done in this paper. As a by-product, our attempt at gap-free law design offers an applicable way to address the responsibility-gap concern in the literature.
2 Formalisation
In this section, we first define a (one-shot current) game to model multiagent systems. Then, we formalise the concepts of law, usefulness, and gap-freeness as discussed in the previous section. After that, we formally define hypergraphs that are used later to solve the law design problems.
In the rest of this paper, by $|S|$ we denote the size of a set $S$; by $\prod F$ and $\bigcup F$ we denote, respectively, the Cartesian product and the union of all sets in a family $F$. For an indexed set $\alpha=\{\alpha_{i}\}_{i\in I}$, by $\mathcal{S}(\alpha)=\{\alpha_{i}\mid i\in I\}$ we denote the support set of $\alpha$ that forgets the order and multiplicity.
Definition 1.
A game is a tuple $(\mathcal{A},\Delta,\mathbb{P})$ such that
  1. 1.
    $\mathcal{A}$ is a nonempty finite set of agents;
  2. 2.
    $\Delta=\{\Delta_{a}\}_{a\in\mathcal{A}}$ is a family of sets of actions, where $\Delta_{a}$ is a finite set of actions available to agent $a$;
  3. 3.
    $\mathbb{P}\subseteq\prod\Delta$ is the prohibition.
In a game $(\mathcal{A},\Delta,\mathbb{P})$, each profile $\delta\in\prod\Delta$ represents an outcome. Instead of utilities of agents [^1] We do not yet consider the incentives of agents in the law design problem, and thus get rid of the utility functions in our model. , we consider a set $\mathbb{P}$ of prohibited outcomes. The purpose of designing laws is to avoid the game ending in the prohibited outcomes. For instance, in the factory example, all factories dumping on the first day represents an outcome where the fish is killed, and the set of prohibited outcomes consists of the three cases where all agents dump on the same day during each three-day cycle. A law is established to avoid the death of the fish, that is, to prevent any of the prohibited outcomes.
Technically, item 2 of Definition 1 allows the action set $\Delta_{a}$ to be empty for any agent $a\in\mathcal{A}$. This is made for mathematical convenience. Moreover, a profile $\delta\in\prod\Delta$ is essentially a set indexed by every agent $a\in\mathcal{A}$ (i.e. $\delta=\{\delta_{a}\}_{a\in\mathcal{A}}$). Correspondingly,
$\mathcal{S}(\delta)=\{\delta_{a}\mid a\in\mathcal{A}\}$ (1)
is the set of actions taken by the agents under a profile $\delta$. Also, the action sets of different agents are not necessarily identical or disjoint. This allows greater flexibility in modelling multiagent systems that simultaneously include multiple agents of the same type who are likely to share the same action space, and agents of different types whose available actions usually differ. More importantly, the laws that set bans on actions, as discussed in Section 1 and formally defined below, will be agent-independent. In other words, it is impossible to forbid one agent from using an action while permitting another agent to use the same action. In this sense, fairness is embedded in our formalisation.
Definition 2.
A law in a game $(\mathcal{A},\Delta,\mathbb{P})$ is an arbitrary set $L\subseteq\bigcup\Delta$ of actions.
For two laws $L,L^{\prime}$ in the same game, we say that $L$ is a reduction of $L^{\prime}$ if $L\subseteq L^{\prime}$. Note that, as stated in Section 1, agents can break a law. However, in a hypothetical situation where all agents obey the law, the agents indeed play a “subgame” where only lawful actions are available. We call such a “subgame” law-imposed game and formalise it below.
Definition 3.
For a law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$, the law-imposed game is the game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$ where
  1. 1.
    $\Delta^{L}=\{\Delta_{a}^{L}\}_{a\in\mathcal{A}}$ such that $\Delta_{a}^{L}=\Delta_{a}\setminus L$ is the set of lawful actions of agent $a$;
  2. 2.
    $\mathbb{P}^{L}=\mathbb{P}\cap\prod\Delta^{L}$ is the law-imposed prohibition.
Informally, for a given game and a given law, in the law-imposed game, the actions are the lawful actions, the profiles are the lawful profiles, and the prohibition consists of the lawful profiles prohibited in the original game.
2.1 Usefulness of Law
Let us now formalise usefulness. Recall that, as discussed in Section 1, a law is useful if the prohibited outcomes never appear when every agent obeys the law. That means none of the lawful profiles are prohibited, which is formalised below.
Definition 4.
A law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$ is called useful if $\mathbb{P}^{L}=\varnothing$ in the law-imposed game.
The next lemma characterises when a law is useful: every prohibited profile consists of at least one banned action by the law. See its formal proof in Appendix A.1.
Lemma 1.
A law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$ is useful if and only if $L\cap\mathcal{S}(\delta)\neq\varnothing$ for each profile $\delta\in\mathbb{P}$.
Next, we consider the minimality of a useful law. Our discussion about it in Section 1 can be formally captured below.
Definition 5.
A law $L$ in a game is minimal-useful if $L$ is useful and no law $L^{\prime}\subsetneq L$ is useful in the same game.
Motivated by the pursuit of minimum constraints on society, if a law is useful but not minimal, then we would reduce it to leave as much freedom (i.e. lawful actions) as possible. This is formally captured as the minimality of useful reduction in the definition below.
Definition 6.
For a useful law $L$ in a game,
  1. 1.
    a useful reduction of $L$ is a useful law $L^{\prime}\subseteq L$;
  2. 2.
    a useful reduction $L^{\prime}$ of $L$ is called minimum if there is no useful reduction $L^{\prime\prime}$ of $L$ such that $|L^{\prime\prime}|<|L^{\prime}|$.
Note that the approach of reducing an existing law, rather than crafting a new one from scratch, aligns with the view that laws should be stable (jefferson1787unstability; raz1979rule), as it ensures that actions previously permitted remain permitted. This is indeed a type of Pareto optimisation. Also, it is not hard to deduce from Lemma 1 that any further reduction of a non-useful law is also non-useful, which is why we consider only the reduction of useful laws here. Moreover, observe that the law which bans all actions (i.e. $L=\bigcup\Delta$) is useful by Lemma 1, and thus crafting a new useful law from scratch is equivalent to getting a useful reduction of the law $L=\bigcup\Delta$. In this sense, minimising an existing useful law encompasses the task of minimum useful law design.
Notably, the law $L=\bigcup\Delta$ may not be reducible while preserving usefulness. For example, consider a matching-pennies game where two agents have the same action space $\{\mathit{head},\mathit{tail}\}$ and the outcomes $(\mathit{head},\mathit{head})$ and $(\mathit{tail},\mathit{tail})$ are prohibited. In this case, the set $\{\mathit{head},\mathit{tail}\}$ is the only useful law by Lemma 1 and thus minimal-useful. In other words, some games may admit no useful law that permits any lawful actions. However, such situations are uncommon in real-world scenarios, where every agent usually has access to a default action that can avoid undesirable outcomes. As a result, a law that permits only default actions will be useful, and any useful reduction of such a law continues to allow those default actions. In the worst-case scenario where no suitable default action exists, a useful initial law can still be constructed by distinguishing between different agents’ actions (i.e. by making their action sets disjoint), albeit at the cost of symmetry. Then, a law that restricts each agent to a single action leading to a fixed non-prohibited outcome (e.g. banning one agent from choosing $\mathit{head}$ and the other from choosing $\mathit{tail}$ in the matching-pennies game) is always useful. This approach mirrors how traffic lights prevent collisions at intersections: when east-west traffic is permitted to proceed, north-south traffic is banned, and vice versa.
2.2 Gap-Freeness of Law
In this subsection, we formalise gap-freeness. As discussed in Section 1, a key feature of gap-freeness is the presence of a principal agent who can individually prevent the prohibited outcomes. In our one-shot game setting, such an ability to prevent is achieved with a “safe” action that guarantees non-prohibited outcomes. We formally capture this as follows.
Definition 7.
For a game $(\mathcal{A},\Delta,\mathbb{P})$ and an agent $a\in\mathcal{A}$, an action $d\in\bigcup\Delta$ is called a safe action of agent $a$ if $d\in\Delta_{a}$ and $\delta_{a}\neq d$ for each profile $\delta\in\mathbb{P}$.
Informally, action $d$ is available to agent $a$, but she does not choose $d$ in any prohibited outcome. In other words, every potential outcome when agent $a$ chooses action $d$ is not prohibited. As a result, agent $a$ can prevent the prohibited outcomes by choosing action $d$. Also note that a safe action $d$ of agent $a$ is not necessarily safe for another agent $b$, even if $d\in\Delta_{a}\cap\Delta_{b}$. This arises from the asymmetry across agents in the prohibition as specified in item 3 of Definition 1.
However, as noted in Section 1, with a law in place, a principal agent needs only a lawful action that prevents the prohibited outcomes when others act lawfully. That is essentially a safe action in the law-imposed game. Next, we capture the legal responsibility of an agent breaking a law and the counterfactual responsibility of a principal agent failing to prevent. In particular, we say that an agent is responsible if a prohibited outcome happens and she is responsible either legally or counterfactually, as defined below.
Definition 8.
In a prohibited profile $\delta\in\mathbb{P}$ of a game $(\mathcal{A},\Delta,\mathbb{P})$, an agent $a\in\mathcal{A}$ is responsible under law $L$ if
  1. 1.
    (legally) $\delta_{a}\in L$, or
  2. 2.
    (counterfactually) $\mathcal{S}(\delta)\cap L=\varnothing$ and a safe action of agent $a$ exists in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$.
Informally, in a prohibited outcome $\delta$, agent $a$ is legally responsible if her action breaks the law (i.e. $\delta_{a}\in L$); if no agent breaks the law (i.e. $\mathcal{S}(\delta)\cap L=\varnothing$) but agent $a$ is a principal agent under the law (i.e. has a safe action in the law-imposed game), then $a$ is counterfactually responsible.
The next lemma characterises when an action is a safe action of an agent under a law: the law should make the action lawful and make each prohibited outcome where the agent takes the action unlawful. See Appendix A.2 for its proof.
Lemma 2.
An action $d\in\bigcup\Delta$ is a safe action of an agent $a\in\mathcal{A}$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$ if and only if $d\in\Delta_{a}^{L}$ and $L\cap\mathcal{S}(\delta)\neq\varnothing$ for each $\delta\in\mathbb{P}$ such that $\delta_{a}=d$.
The next definition formalises the gap-freeness property.
Definition 9.
A law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$ is gap-free if there is at least one responsible agent in each profile $\delta\in\mathbb{P}$.
Observe that, by Lemma 1 and statement (1), if a law is useful, then there is at least one legally responsible agent in each prohibited profile. Thus, a useful law is also gap-free. Conversely, if a law is not useful, then there is at least one prohibited profile where no agent is legally responsible. To make such a law gap-free, there should be a principal agent who bears counterfactual responsibility in those lawful but prohibited profiles. This intuition is formally captured by the next lemma. See Appendix A.3 for its formal proof.
Lemma 3.
A law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$ is gap-free if and only if $L$ is useful or there is an agent $a\in\mathcal{A}$ and a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$.
Now, we formalise the minimality of gap-freeness and gap-free reduction in a manner analogous to Section 2.1.
Definition 10.
A law $L$ in a game is minimal-gap-free if $L$ is gap-free and no law $L^{\prime}\subsetneq L$ is gap-free in the same game.
Definition 11.
For a gap-free law $L$ in a game,
  1. 1.
    a gap-free reduction of $L$ is a gap-free law $L^{\prime}\subseteq L$;
  2. 2.
    a gap-free reduction $L^{\prime}$ of $L$ is called minimum if there is no gap-free reduction $L^{\prime\prime}$ of $L$ such that $|L^{\prime\prime}|<|L^{\prime}|$.
Recall that the law $\bigcup\Delta$ is useful and thus gap-free by Lemma 3. Hence, designing a gap-free law is equivalent to finding a gap-free reduction of the law $\bigcup\Delta$. Accordingly, minimising an existing gap-free law encompasses the task of minimum gap-free law design.
2.3 Hypergraphs with Fixed Rank
Now, let us introduce the hypergraphs that are used later to resolve the law design problems. Unlike standard graphs where each edge connects exactly two vertices, a hypergraph allows each edge (a.k.a. hyperedge) to connect any positive number of vertices. In particular, we consider hypergraphs where each edge contains at most $k$ vertices for any fixed parameter $k\geq 1$, as formalised in the next definition.
Definition 12.
For any integer $k\geq 1$, a rank-$k$ hypergraph (abbr. $k$-graph) is a tuple $(V,E)$ such that $V$ is a finite set of vertices and $E$ is a set of (hyper)edges where $e\subseteq V$ and $1\leq|e|\leq k$ for each edge $e\in E$.
Note that an edge is a set of vertices. A vertex is said to cover an edge if the edge includes the vertex. A vertex cover of a $k$-graph is a set of vertices that intersects with every edge in the $k$-graph, as formalised below.
Definition 13.
For a $k$-graph $(V,E)$,
  1. 1.
    a set $C$ is called a vertex cover if $C\subseteq V$ and $C\cap e\neq\varnothing$ for each edge $e\in E$;
  2. 2.
    a vertex cover $C$ is called minimal if there is no vertex cover $C^{\prime}\subsetneq C$;
  3. 3.
    a vertex cover $C$ is called minimum if there is no vertex cover $C^{\prime}$ such that $|C^{\prime}|<|C|$.
We consider the following problems about vertex cover, which are collectively referred to as $\mathbf{VC}$ problems:
  • $\mathbf{IsVC}$: to verify if a set is a vertex cover of a $k$-graph.
  • $\mathbf{IsMiniVC}$: to verify if a set is a minimal vertex cover of a $k$-graph.
  • $\mathbf{MinVC}$: to find a minimum vertex cover of a $k$-graph.
These problems are extensively explored in the literature. In particular, $\mathbf{IsVC}$ and $\mathbf{IsMiniVC}$ can both be solved efficiently (i.e. in polynomial time). In contrast, $\mathbf{MinVC}$ is NP-hard (garey1979computers), which means no efficient algorithm is believed to exist for this problem. As a compromise, approximation algorithms were developed to efficiently find good enough, though not necessarily optimal, solutions for those hard problems. An algorithm approximates $\mathbf{MinVC}$ within factor $t$ if, for any $k$-graph, the size of the vertex cover it finds is at most $t$ times the size of the minimum vertex cover. It is shown that $\mathbf{MinVC}$ can be effectively approximated within factor $k$, using greedy algorithms (bar1981linear; hall1986fast) or linear programming relaxation (hochbaum1982approximation). However, it is hard to do better than this (holmerin2002improved; dinur2005new). In what follows, we will rely on the following algorithms and theorem.
Algorithms.
While solving the law design problems, the following efficient $\mathbf{VC}$ algorithms serve as gadgets:
  • $\mathtt{IsVC}$: an algorithm for $\mathbf{IsVC}$.
  • $\mathtt{IsMiniVC}$: an algorithm for $\mathbf{IsMiniVC}$.
  • $\mathtt{AppMinVC}$: a $k$-approximation algorithm for $\mathbf{MinVC}$.
Theorem 1 (khot2008vertex, khot2008vertex).
$\mathbf{MinVC}$ is NP-hard to approximate within factor $k-\varepsilon$ for any $\varepsilon\!>\!0$ when $k\!\geq\!2$.222khot2008vertex consider a subset of $k$-graphs where each edge has exactly $k$ vertices. The theorem holds under the Unique Game Conjecture (khot2002power).
3 Useful Law Design
To continue the discussion in Section 2.1, we consider the next three problems about useful law design, which are collectively referred to as $\mathbf{UL}$ problems:
  • $\mathbf{IsUL}$: to verify if a set is a useful law in a game.
  • $\mathbf{IsMiniUL}$: to verify if a set is a minimal-useful law in a game.
  • $\mathbf{MinUR}$: to find a minimum useful reduction of a useful law in a game.
In particular, we establish an equivalence between the $\mathbf{VC}$ problems and the $\mathbf{UL}$ problems by providing two-way polynomial-time reductions. On top of this, we illustrate a way to solve the $\mathbf{UL}$ problems using the $\mathbf{VC}$ algorithms.
3.1 Reducing Vertex Cover to Useful Law
In this subsection, we show that any instance of a $\mathbf{VC}$ problem can be reduced to an instance of a $\mathbf{UL}$ problem in polynomial time. This demonstrates that the $\mathbf{UL}$ problems are at least as hard as the $\mathbf{VC}$ problems. The polynomial-time reduction is formally captured below.
Definition 14.
For any $k$-graph $(V,E)$, let $\mathcal{G}_{(k,V,E)}$ be the game $(\mathcal{A}_{(k,V,E)},\Delta_{(k,V,E)},\mathbb{P}_{(k,V,E)})$ such that
  1. 1.
    $\mathcal{A}_{(k,V,E)}=[k]$ where $k]=\{i\in\mathbb{N}\mid 1\leq i\leq k\}$;
  2. 2.
    $\Delta_{(k,V,E)}=\{\Delta_{i}\}_{i\in[k]}$ where $\Delta_{i}=V$ for each $i\in[k]$;
  3. 3.
[…]
Note that, in the above definition, $k\geq 1$ and set $V$ is finite by Definition 12. Thus, $\mathcal{G}_{(k,V,E)}$ is well-defined by Definition 1. Informally, $\mathcal{G}_{(k,V,E)}$ is a game of $k$ agents with the same action space $V$. Each vertex in set $V$ is an action. Each edge $e\in E$ corresponds to a prohibited profile $\delta^{e}$ where every vertex in edge $e$ is taken as an action $\delta^{e}_{i}$ of some agent $i$. Then, by Lemma 1 and item 1 of Definition 13, it is easy to verify the next theorem. See Appendix B.1 for its proof.
Theorem 2.
A set $C$ is a vertex cover of a $k$-graph $(V,E)$ if and only if $C$ is a useful law in the game $\mathcal{G}_{(k,V,E)}$.
Note that, due to the consistency of minimality in item 2 of Definition 13 and Definition 5, Theorem 2 indeed implies reductions from $\mathbf{IsVC}$ and $\mathbf{IsMiniVC}$ to $\mathbf{IsUL}$ and $\mathbf{IsMiniUL}$, respectively. On the other hand, the set $V=\bigcup\Delta_{(k,V,E)}$ is a useful law in the game $\mathcal{G}_{(k,V,E)}$ by Definition 14 and Lemma 1. Then, every useful law in the game $\mathcal{G}_{(k,V,E)}$ is a useful reduction of law $V$ by item 1 of Definition 6. Thus, the next corollary follows from Theorem 2.
Corollary 1.
A set $C$ is a vertex cover of a $k$-graph $(V,E)$ if and only if $C$ is a useful reduction of law $V$ in game $\mathcal{G}_{(k,V,E)}$.
Note that, due to the consistency of minimality in item 3 of Definition 13 and item 2 of Definition 6, Corollary 1 implies a reduction from $\mathbf{MinVC}$ to $\mathbf{MinUR}$. This, together with Theorem 1, further implies an inapproximability result of the $\mathbf{MinUR}$ problem as stated in the next theorem. See Appendix B.2 for its proof.
Theorem 3.
$\mathbf{MinUR}$ in a game $(\mathcal{A},\Delta,\mathbb{P})$ is NP-hard to approximate within factor $|\mathcal{A}|-\varepsilon$ for any $\varepsilon\!>\!0$ when $|\mathcal{A}|\!\geq\!2$.
3.2 Reducing Useful Law to Vertex Cover
In this subsection, we show that any instance of a $\mathbf{UL}$ problem can be polynomially reduced to an instance of a $\mathbf{VC}$ problem. This demonstrates that the $\mathbf{UL}$ problems are no harder than the $\mathbf{VC}$ problems and suggests how $\mathbf{VC}$ algorithms can be used to solve the $\mathbf{UL}$ problems.
[…]
Theorem 4.
A set $L$ is a useful law in a game $(\mathcal{A},\Delta,\mathbb{P})$ if and only if $L$ is a vertex cover of the graph $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$.
Opposite to Theorem 2, Theorem 4 implies reductions from $\mathbf{IsUL}$ and $\mathbf{IsMiniUL}$ to $\mathbf{IsVC}$ and $\mathbf{IsMiniVC}$, respectively. The corresponding algorithms $\mathtt{IsUL}$ and $\mathtt{IsMiniUL}$ call $\mathtt{IsVC}$ and $\mathtt{IsMiniVC}$, respectively, as shown in Algorithm 1 of Appendix B.6.
Next, we consider the $\mathbf{MinUR}$ problem. Note that, by Theorem 4, reducing a useful law corresponds to finding a smaller vertex cover within an existing one. Meanwhile, the smaller vertex cover intersects every edge at the vertices in the original cover. Given this observation, the smaller vertex cover can be regarded as a vertex cover in a subgraph induced by the original cover, which is formalised as follows.
Definition 15.
For a vertex cover $C$ of a $k$-graph $(V,E)$, the induced subgraph is the $k$-graph $(C,E^{C})$ such that $E^{C}=\{C\cap e\mid e\in E\}$.
The subgraph $(C,E^{C})$ is well-defined because $C\cap e\neq\varnothing$ and $|C\cap e|\leq|e|\leq k$ for each edge $e\in E$ as $C$ is a vertex cover in the $k$-graph $(V,E)$. Moreover, observe that a vertex cover of the subgraph $(C,E^{C})$ is also a vertex cover of the graph $(V,E)$. Then, Theorem 4 implies the next theorem. See Appendix B.5 for its formal proof.
Theorem 5.
For a useful law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$, law $L^{\prime}$ is a useful reduction of $L$ if and only if $L^{\prime}$ is a vertex cover of the induced subgraph $(L,\mathcal{S}(\mathbb{P})^{L})$.
In contrast with Corollary 1, Theorem 5 implies a reduction from $\mathbf{MinUR}$ to $\mathbf{MinVC}$ and an algorithm $\mathtt{AppMinUR}$ that calls $\mathtt{AppMinVC}$ (see Algorithm 1 of Appendix B.6). Note that the graph $(L,\mathcal{S}(\mathbb{P})^{L})$ in Theorem 5 is an $|\mathcal{A}|$-graph, on which $\mathtt{AppMinVC}$ is an $|\mathcal{A}|$-approximation. This makes $\mathtt{AppMinUR}$ an $|\mathcal{A}|$-approximation of $\mathbf{MinUR}$, where $|\mathcal{A}|$ is the number of agents in the input game (see Appendix B.7 for a discussion). It means $\mathtt{AppMinUR}$ achieves a nearly optimal approximation factor by Theorem 3.
4 Gap-Free Law Design
Following Section 2.2, in this section, we consider the next three problems about gap-free law design, which are collectively referred to as $\mathbf{GFL}$ problems:
  • $\mathbf{IsGFL}$: to verify if a set is a gap-free law in a game.
  • $\mathbf{IsMiniGFL}$: to verify if a set is a minimal-gap-free law in a game.
  • $\mathbf{MinGFR}$: to find a minimum gap-free reduction of a gap-free law in a game.
We first show the hardness of the $\mathbf{GFL}$ problems by reductions from the $\mathbf{UL}$ problems. Then, we provide a way to solve the $\mathbf{GFL}$ problems again using the $\mathbf{VC}$ algorithms.
4.1 Reducing Useful Law to Gap-Free Law
In this subsection, we show that any instance of a $\mathbf{UL}$ problem can be polynomially reduced to an instance of a $\mathbf{GFL}$ problem. Specifically, for any game $(\mathcal{A},\Delta,\mathbb{P})$, we construct a game $(\bar{\mathcal{A}},\bar{\Delta},\bar{\mathbb{P}})$ as illstrated in the next definition.
Definition 16.
For a game $(\mathcal{A},\Delta,\mathbb{P})$, an agent $\gamma\notin\mathcal{A}$, and two distinct actions $p,n\notin\bigcup\Delta$, let the game $(\bar{\mathcal{A}},\bar{\Delta},\bar{\mathbb{P}})$ be:
  1. 1.
    $\bar{\mathcal{A}}=\mathcal{A}\cup\{\gamma\}$;
  2. 2.
    $\bar{\Delta}=\{\bar{\Delta}_{a}\}_{a\in\bar{\mathcal{A}}}$ where $\bar{\Delta}_{a}=\begin{cases}\Delta_{a}\cup\{n\},&\!\text{if }a\in\mathcal{A};\\ \{p,n\},&\!\text{if }a=\gamma;\end{cases}$
  3. 3.
    $\bar{\mathbb{P}}=\bar{\mathbb{P}}_{1}\cup\bar{\mathbb{P}}_{2}\cup\bar{\mathbb{P}}_{3}$ where
    • $\bar{\mathbb{P}}_{1}=\{\bar{\delta}\mid\bar{\delta}_{\gamma}=p,(\exists\delta\in\mathbb{P},\forall a\in\mathcal{A},\bar{\delta}_{a}=\delta_{a})\}$;
    • $\bar{\mathbb{P}}_{2}=\{\bar{\delta}\mid\exists a\in\mathcal{A}\,(\bar{\delta}_{a}\in\Delta_{a},\forall b\in\bar{\mathcal{A}}\setminus\{a\},\bar{\delta}_{b}=n)\}$;
    • $\bar{\mathbb{P}}_{3}=\{\bar{\delta}\mid\forall a\in\bar{\mathcal{A}},\bar{\delta}_{a}=n\}$.
Observe that the new agent $\gamma$ takes action $n$ in each profile in sets $\bar{\mathbb{P}}_{2},\bar{\mathbb{P}}_{3}$. Meanwhile, by Lemma 1 and the definition of $\bar{\mathbb{P}}_{1}$, a useful law $L$ in the game $(\mathcal{A},\Delta,\mathbb{P})$ intersects with each profile in set $\bar{\mathbb{P}}_{1}$. Then, law $L$ makes $p$ a safe action of agent $\gamma$ in the game $(\bar{\mathcal{A}},\bar{\Delta}^{L},\bar{\mathbb{P}}^{L})$ by Lemma 2. Thus, law $L$ is a gap-free law in the game $(\bar{\mathcal{A}},\bar{\Delta},\bar{\mathbb{P}})$ by Lemma 3. Theorem 6 below formalises the above observation and establishes its converse as well. See Appendix C.1 for its formal proof.
Theorem 6.
A set $L\subseteq\bigcup\Delta$ is a useful law in a game $(\mathcal{A},\Delta,\mathbb{P})$ if and only if $L$ is a gap-free law in the game $(\bar{\mathcal{A}},\bar{\Delta},\bar{\mathbb{P}})$.
Note that, due to the consistency of minimality in Definition 5 and Definition 10, Theorem 6 indeed implies reductions from $\mathbf{IsUL}$ and $\mathbf{IsMiniUL}$ to $\mathbf{IsGFL}$ and $\mathbf{IsMiniGFL}$, respectively. Moreover, when considering a reduction $L^{\prime}$ of a useful law $L$ in the game $(\mathcal{A},\Delta,\mathbb{P})$, it is guaranteed that $L^{\prime}\subseteq L\subseteq\bigcup\Delta$. Then, by item 1 of Definition 6 and item 1 of Definition 11, Theorem 6 further implies the next corollary.
Corollary 2.
For a useful law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$, a set $L^{\prime}$ is a useful reduction of $L$ in the game $(\mathcal{A},\Delta,\mathbb{P})$ if and only if $L^{\prime}$ is a gap-free reduction of $L$ in the game $(\bar{\mathcal{A}},\bar{\Delta},\bar{\mathbb{P}})$.
Given the consistency of minimality in item 2 of Definition 6 and item 2 of Definition 11, Corollary 2 implies a reduction from $\mathbf{MinUR}$ to $\mathbf{MinGFR}$. This, together with Theorem 3 and the extra agent $\gamma$ in Definition 16, further implies an inapproximability result of $\mathbf{MinGFR}$ as stated in the next theorem. See Appendix C.2 for its proof.
Theorem 7.
$\mathbf{MinGFR}$ in a game $(\mathcal{A},\Delta,\mathbb{P})$ is NP-hard to approximate within factor $|\mathcal{A}|-1-\varepsilon$ for any $\varepsilon\!>\!0$ when $|\mathcal{A}|\!\geq\!3$.
4.2 Reducing Gap-Free Law to Vertex Cover
In this subsection, we reduce the $\mathbf{GFL}$ problems to the $\mathbf{VC}$ problems. By this means, we show a way to address the $\mathbf{GFL}$ problems using the $\mathbf{VC}$ algorithms.
Recall Lemma 3 that the gap-freeness of a law corresponds to the usefulness of the law and the existence of a safe action in the law-imposed game. Section 3.2 demonstrated how the $\mathbf{UL}$ problems can be addressed using the $\mathbf{VC}$ algorithms. Now, we discuss how a “safe action in a law-imposed game” is captured in the $\mathbf{VC}$ context. Note that not every action can become a safe action in a law-imposed game. If it can, we say the action is safable.
Definition 17.
For a game $(\mathcal{A},\Delta,\mathbb{P})$, an action $d\in\bigcup\Delta$ is safable if there is a law $L$ and an agent $a$ such that $d$ is a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$.
The next lemma characterises when an action is safable: every agent taking it does not lead to a prohibited outcome. See Appendix C.3 for its formal proof.
Lemma 4.
For a game $(\mathcal{A},\Delta,\mathbb{P})$, an action $d\in\bigcup\Delta$ is safable if and only if $\mathcal{S}(\delta)\neq\{d\}$ for each profile $\delta\in\mathbb{P}$.
Next, let us recall Lemma 2, which characterises when an action $d$ is a safe action of agent $a$ in a law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$. Note that the second half of Lemma 2 implies that $L\subseteq\bigcup\Delta\setminus\{d\}$ and $L\cap(\mathcal{S}(\delta)\setminus\{d\})\neq\varnothing$ for each profile $\delta\in\mathbb{P}$ such that $\delta_{a}=d$. This, by item 1 of Definition 13, implies that $L$ is a vertex cover in the graph defined below.
Definition 18.
[…]
  1. 1.
    $\mathcal{V}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}=(\bigcup\Delta)\setminus\{d\}$;
  2. 2.
    $\mathcal{E}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}=\{\mathcal{S}(\delta)\setminus\{d\}\mid\delta\in\mathbb{P},\delta_{a}=d\}$.
Note that the $|\mathcal{A}|$-graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$ is well-defined because $1\!\leq\!|\mathcal{S}(\delta)\setminus\{d\}|\!<\!|\mathcal{S}(\delta)|\!\leq\!|\mathcal{A}|$ by Lemma 4 and statement (1). Following the above observation, the next lemma bridges “a safe action in a law-imposed game” with the $\mathbf{VC}$ problems. See Appendix C.4 for its formal proof.
Lemma 5.
For a law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$ and an agent $a\in\mathcal{A}$, an action $d\in\bigcup\Delta$ is a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$ if and only if $\in\Delta_{a}^{L}, dd$ is safable in the game $(\mathcal{A},\Delta,\mathbb{P})$, and $L$ is a vertex cover in the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$.
Observe that Lemma 5, Theorem 4, and Lemma 3 imply Theorem 8 below, which further implies a Cook reduction from $\mathbf{IsGFL}$ to $\mathbf{IsVC}$ and a corresponding algorithm $\mathtt{IsGFL}$ that iteratively calls $\mathtt{IsVC}$ for polynomial times, as illustrated in Algorithm 2 of Appendix C.8.
Theorem 8.
A law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$ is gap-free if and only if at least one of the following statements is true:
  1. 1.
    $L$ is a vertex cover in the graph $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$;
  2. 2.
    there is an agent $a\in\mathcal{A}$ and a safable action $d\in\Delta_{a}^{L}$ such that $L$ is a vertex cover in the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$.
Note that, by Definition 10 and item 1 of Definition 11, a gap-free law is not minimal if there is a “strict” gap-free reduction. By Lemma 3, such a reduction retains the gap-freeness in three possible approaches: maintaining usefulness, maintaining a safe action under the original law, or introducing a new safe action. Moreover, we can use an induced subgraph in Definition 15 to capture a reduction of an existing law. Following the above hints, we can get the next two theorems. See Appendices C.5 and C.6 for their proofs.
Theorem 9.
A gap-free law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$ is minimal if and only if all of the following statements are true:
  1. 1.
    if $L$ is a vertex cover of the graph $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$, then $L$ is a minimal vertex cover in this graph;
  2. 2.
    for each agent $a\in\mathcal{A}$ and each safable action $d\in\Delta_{a}^{L}$, if $L$ is a vertex cover in the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$, then $L$ is a minimal vertex cover in this graph;
  3. 3.
    for each agent $a\in\mathcal{A}$ and each safable action $d\!\in\!\Delta_{a}\!\cap\!L$, the set $L\setminus\{d\}$ is not a vertex cover in graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$.
Theorem 10.
For a gap-free law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$, a law $L^{\prime}$ is a gap-free reduction of $L$ if and only if at least one of the following statements is true:
  1. 1.
    $L$ is a vertex cover of the graph $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$ and $L^{\prime}$ is a vertex cover in the subgraph $(L,\mathcal{S}(\mathbb{P})^{L})$;
  2. 2.
    there is an agent $a\in\mathcal{A}$ and a safable action $d\in\Delta_{a}^{L}$ such that $L$ is a vertex cover in graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$ and $L^{\prime}$ is a vertex cover in the subgraph $(L,(\mathcal{E}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})})^{L})$;
  3. 3.
    there is an $a\!\in\!\mathcal{A}$ and a safable action $d\!\in\!\Delta_{a}\!\cap\!L$ such that $L\!\setminus\!\{d\}$ is a vertex cover in graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$ and $L^{\prime}$ is a vertex cover in the subgraph $(L\!\setminus\!\{d\},(\mathcal{E}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})})^{L\setminus\{d\}})$.
The above two theorems imply Cook reductions from $\mathbf{IsMiniGFL}$ and $\mathbf{MinGFR}$ to the $\mathbf{VC}$ problems. The corresponding algorithms $\mathtt{IsMiniGFL}$ and $\mathtt{AppMinGFR}$ call the $\mathbf{VC}$ algorithms polynomial times (see Algorithm 2 of Appendix C.8). Note that all graphs in Theorem 10 are $|\mathcal{A}|$-graphs, on which $\mathtt{AppMinVC}$ is an $|\mathcal{A}|$-approximation. This makes $\mathtt{AppMinGFR}$ an $|\mathcal{A}|$-approximation of $\mathbf{MinGFR}$, where $|\mathcal{A}|$ is the number of agents in the input game (see Appendix C.7 for a discussion).
5 Conclusion
For the usefulness and gap-freeness of law, we studied the corresponding law design problems by relating them to vertex cover problems in hypergraphs. We proved that the task of minimising a law while keeping its usefulness or gap-freeness is NP-hard even to approximate. We also proposed reductions from law design problems to vertex cover problems, which imply law-design algorithms that make polynomial-time calls to the vertex cover algorithms.
As an ending discussion, note that we don’t consider the weight of actions in law minimisation. However, this can be done with no extra effort because the vertex cover algorithms in the literature apply to weighted vertices.
Technical Appendix
Appendix A Supplementary to Section 2
A.1 Proof of Lemma 1
Lemma 1. A law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$ is useful if and only if $L\cap\mathcal{S}(\delta)\neq\varnothing$ for each profile $\delta\in\mathbb{P}$.
Proof.
By Definition 4, it suffices to show that $\mathbb{P}^{L}=\varnothing$ in the law-imposed game if and only if $L\cap\mathcal{S}(\delta)\neq\varnothing$ for each profile $\delta\in\mathbb{P}$.
($\Rightarrow$) Suppose there is a profile $\delta\in\mathbb{P}$ such that $L\cap\mathcal{S}(\delta)=\varnothing$. Then, $\delta_{a}\notin L$ for each agent $a\in\mathcal{A}$ by statement (1). Thus, $\delta\in\prod\Delta^{L}$ by item 1 of Definition 3. Hence, $\delta\in\mathbb{P}^{L}$ by item 2 of Definition 3 and the case assumption $\delta\in\mathbb{P}$. Therefore, $\mathbb{P}^{L}\neq\varnothing$.
($\Leftarrow$) Suppose $\mathbb{P}^{L}\neq\varnothing$. Then, there is a profile $\delta\in\mathbb{P}^{L}$. Thus, $\delta\in\prod\Delta^{L}$ by item 2 of Definition 3. Hence, $\delta_{a}\notin L$ for each agent $a\in\mathcal{A}$ by item 1 of Definition 3. Therefore, $L\cap\mathcal{S}(\delta)=\varnothing$ by statement (1). ∎
A.2 Proof of Lemma 2
Lemma 2. An action $d\in\bigcup\Delta$ is a safe action of an agent $a\in\mathcal{A}$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$ if and only if $d\in\Delta_{a}^{L}$ and $L\cap\mathcal{S}(\delta)\neq\varnothing$ for each $\delta\in\mathbb{P}$ such that $\delta_{a}=d$.
Proof.
By Definition 7 and Definition 3, action $d$ is a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$ if and only if $d\in\Delta_{a}^{L}$ and $\delta_{a}\neq d$ for each profile $\delta\in\mathbb{P}^{L}$. Then, to prove the statement of the lemma, it suffices to prove the equivalence between the next two statements: (I) $\delta_{a}\neq d$for each profile $\delta\in\mathbb{P}^{L}$; (II) $L\cap\mathcal{S}(\delta)\neq\varnothingfor$ each profile $\delta\in\mathbb{P}$ such that $\delta_{a}=d$.
(I) $\Rightarrow$ (II): Suppose there exists a profile $\delta\in\mathbb{P}$ such that $\delta_{a}=d$ and $L\cap\mathcal{S}(\delta)=\varnothing$. Then, $\delta_{b}\notin L$ for each agent $b\in\!\mathcal{A}$ by statement (1). Thus, $\delta$ is such that $\delta\in\mathbb{P}^{L}$ and $\delta_{a}=d$ by Definition 3 and the assumptions $\delta\in\mathbb{P}$ and $\delta_{a}=d$.
(II) $\Rightarrow$ (I): Suppose there is a profile $\delta\in\mathbb{P}^{L}$ such that $\delta_{a}=d$. Then, $\delta\in\mathbb{P}$ and $\delta_{b}\notin L$ for each agent $b\in\mathcal{A}$ by Definition 3. Thus, $\delta$ is such that $\delta\in\mathbb{P}$ and $L\cap\mathcal{S}(\delta)=\varnothing$ by statement (1). ∎
A.3 Proof of Lemma 3
Lemma 3. A law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$ is gap-free if and only if $L$ is useful or there is an agent $a\in\mathcal{A}$ and a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$.
Proof.
($\Rightarrow$) Suppose $L$ is gap-free but not useful. Then, $\mathbb{P}^{L}\neq\varnothing$ by Lemma 1. Consider an arbitrary profile $\delta\in\mathbb{P}^{L}$. Then, $\delta\in\mathbb{P}$ and
$\delta_{b}\notin L$ for each agent $b\in\mathcal{A}$ (2)
by Definition 3. Thus, there is an agent $a\in\mathcal{A}$ responsible for $\delta$ by Definition 9 and the assumption $L$ is gap-free. Then, there is a safe action $d$ of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$ by Definition 8 and statement (2).
($\Leftarrow$) Suppose $L$ is not gap-free. Then, by Definition 9,
there is a profile $\delta\in\mathbb{P}$ where no agent is responsible. (3)
Consider an arbitrary agent $a\in\mathcal{A}$. Then, $\delta_{a}\notin L$ by item 1 of Definition 8 and statement (3). Thus, by statement (1), Definition 3, and the arbitrariness of $a\!\in\!\mathcal{A}$,
$\mathcal{S}(\delta)\cap L=\varnothing$ (4)
and $\delta\in\mathbb{P}\cap\prod\Delta^{L}=\mathbb{P}^{L}$. The latter implies $\mathbb{P}^{L}\neq\varnothing$. Hence, law $L$ is not useful by Lemma 1. Meanwhile, by statements (3), (4), and item 2 of Definition 8, there is no safe action of (the arbitrary) agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$. ∎
Appendix B Supplementary to Section 3
B.1 Proof of Theorem 2
To increase the readability of the proof of Theorem 2, let us first prove the next lemma.
Lemma 6.
$\mathcal{S}(\delta^{e})=e$ for each profile $\delta^{e}\in\mathbb{P}_{(k,V,E)}$.
Proof.
($\subseteq$) Consider an arbitrary action $d\in\mathcal{S}(\delta^{e})$. Then, $d=\delta^{e}_{i}$ for some agent $i\in[k]$ by statement (1) and item 1 of Definition 14. Thus, $d\in e$ by item 3 of Definition 14.
($\supseteq$) Consider an arbitrary vertex $v\in e$. Suppose $v$ is the $j^{\text{th}}$ item in the predefined order of set $e$ in item 3 of Definition 14. Then,
$1\leq j\leq|e|.$ (5)
Note that $(V,E)$ is a $k$-graph by Definition 14. Then, $|e|\leq k$ by Definition 12. Thus, $1\leq j\leq|e|\leq k$ by statement (5). Hence, one of the next cases must be true:
Case 1: $j=1$. Then, |e|mod|e|+1=j|e|\ $\mathrm${mod}\ |e|+1=j. Thus, $\delta^{e}_{|e|}=v$ by item 3 of Definition 14. Then, $v\!\in\!\mathcal{S}(\delta^{e})$ by statement (1).
Case 2: $1<j\leq|e|$. Then, (j1)mod|e|+1=j(j-1)\ $\mathrm${mod}\ |e|+1=j. Thus, $\delta^{e}_{j-1}=v$ by item 3 of Definition 14. Then, $v\!\in\!\mathcal{S}(\delta^{e})$ by statement (1). ∎
Lemma 6 above implies that $L\cap e\neq\varnothing$ is true if and only if $L\cap\mathcal{S}(\delta^{e})\neq\varnothing$. This intuitively shows that a vertex set $L$ covers an edge $e$ if and only if the prohibited profile $\delta^{e}$ is unlawful under law $L$. Next, we prove Theorem 2.
Theorem 2. A set $C$ is a vertex cover of a $k$-graph $(V,E)$ if and only if $C$ is a useful law in the game $\mathcal{G}_{(k,V,E)}$.
Proof.
Note that, by item 2 of Definition 14,
$\bigcup\Delta_{(k,V,E)}=V.$ (6)
($\Rightarrow$) Suppose set $C$ is a vertex cover of the graph $(V,E)$. Then, $C\subseteq V$ by item 1 of Definition 13. Thus, set $C$ is a law in the game $\mathcal{G}_{(k,V,E)}$ by Definition 2 and statement (6). Hence, it suffices to show that $C\cap\mathcal{S}(\delta^{e})\neq\varnothing$ for each profile $\delta^{e}\in\mathbb{P}_{(k,V,E)}$ by Lemma 1.
Suppose the opposite. Then, a profile $\delta^{e}\in\mathbb{P}_{(k,V,E)}$ such that $C\cap\mathcal{S}(\delta^{e})=\varnothing$ exists. Thus, $C\cap e=\varnothing$ by Lemma 6. Hence, $C$ is not a vertex cover of the graph $(V,E)$ by item 1 of Definition 13, which contradicts the ($\Rightarrow$) part assumption.
($\Leftarrow$) Suppose $C$ is a useful law in the game $\mathcal{G}_{(k,V,E)}$. Then, by Definition 2 and statement (6),
$C\subseteq V$ (7)
and $C\cap\mathcal{S}(\delta^{e})\neq\varnothing$ for each profile $\delta^{e}\in\mathbb{P}_{(k,V,E)}$ by Lemma 1. Thus, $C\cap e\neq\varnothing$ for each edge $e\in E$ by item 3 of Definition 14 and Lemma 6. Hence, $C$ is vertex cover of the graph $(V,E)$ by statement (7) and item 1 of Definition 13. ∎
B.2 Proof of Theorem 3
Theorem 3. $\mathbf{MinUR}$ in a game $(\mathcal{A},\Delta,\mathbb{P})$ is NP-hard to approximate within factor $|\mathcal{A}|-\varepsilon$ for any $\varepsilon\!>\!0$ when $|\mathcal{A}|\!\geq\!2$.
Proof.
Suppose the opposite. Then, there is an $\varepsilon>0$ and an algorithm $\mathtt{Alg}$ that approximates $\mathbf{MinUR}$ with factor $|\mathcal{A}|-\varepsilon$ for any game $(\mathcal{A},\Delta,\mathbb{P})$ where $|\mathcal{A}|\geq 2$.
Consider an arbitrary $k$-graph $(V,E)$ where $k\geq 2$. We consider the next two steps:
  1. 1.
    compute the game $\mathcal{G}_{(k,V,E)}$ corresponding to the graph $(V,E)$ by Definition 14;
  2. 2.
    use the algorithm $\mathtt{Alg}$ to get an approximated useful reduction $L$ of law $V$ in the game $\mathcal{G}_{(k,V,E)}$.
By step 2 above and Corollary 1, set $L$ is a vertex cover of the graph $(V,E)$.
Note that the minimum useful reduction of law $V$ in the game $\mathcal{G}_{(k,V,E)}$ is the minimum vertex cover of the graph $(V,E)$ by Corollary 1. Meanwhile, the size of set $L$ is at most $(k-\varepsilon)$ times the size of the minimum useful reduction of law $V$ by the assumption about the approximation factor of $\mathtt{Alg}$. Thus, the size of set $L$ is at most $(k-\varepsilon)$ times the size of the minimum vertex cover of the graph $(V,E)$. Hence, by the arbitrariness of the $k$-graph $(V,E)$, the above two steps form an algorithm that approximates $\mathbf{MinVC}$ with factor $k-\varepsilon$, which contradicts Theorem 1. ∎
B.3 the $|\mathcal{A}|$-graph $(\bigcup\Delta,\mathcal{S}(\delta))$
As a standard notation in mathematics,
$\mathcal{S}(\mathbb{P})=\{\mathcal{S}(\delta)\mid\delta\in\mathbb{P}\}.$ (8)
Lemma 7.
For a game $(\mathcal{A},\Delta,\mathbb{P})$, the tuple $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$ is an $|\mathcal{A}|$-graph.
Proof.
Note that, by items 1 and 2 of Definition 1, the sets $\mathcal{A}$ and $\Delta_{a}$ for each $a\in\mathcal{A}$ are all finite. Then,
the set $\bigcup\Delta$ is finite. (9)
At the same time, $\mathcal{A}\neq\varnothing$ by item 1 of Definition 1. Then, $1\leq|\mathcal{S}(\delta)|\leq|\mathcal{A}|$ for each profile $\delta\in\mathbb{P}$ by item 3 of Definition 1 and statement (1). Thus, the statement of the lemma is true by Definition 12 and statement (9). ∎
B.4 Proof of Theorem 4
Theorem 4. A set $L$ is a useful law in a game $(\mathcal{A},\Delta,\mathbb{P})$ if and only if $L$ is a vertex cover of the graph $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$.
Proof.
($\Rightarrow$) Suppose that the set $L$ is a useful law in the game $(\mathcal{A},\Delta,\mathbb{P})$. Then, $L\subseteq\bigcup\Delta$ and $\mathcal{S}(\delta)\cap L\neq\varnothing$ for each profile $\delta\in\mathbb{P}$ by Definition 2 and Lemma 1. Thus, $L$ is a vertex cover of the graph $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$ by item 1 of Definition 13 and statement (8).
($\Leftarrow$) Suppose that the set $L$ is a vertex cover of the graph $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$. Then, $L\subseteq\bigcup\Delta$ and $\mathcal{S}(\delta)\cap L\neq\varnothing$ for each profile $\delta\in\mathbb{P}$ by item 1 of Definition 13 and statement (8). Thus, $L$ is a useful law in the game $(\mathcal{A},\Delta,\mathbb{P})$ by Definition 2 and Lemma 1. ∎
B.5 Proof of Theorem 5
Theorem 5. For a useful law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$, law $L^{\prime}$ is a useful reduction of $L$ if and only if $L^{\prime}$ is a vertex cover of the induce subgraph $(L,\mathcal{S}(\mathbb{P})^{L})$.
Proof.
($\Rightarrow$) Suppose $L^{\prime}$ is not a vertex cover of the graph $(L,\mathcal{S}(\mathbb{P})^{L})$. Then, by items 1 of Definition 13, one of the next cases must be true:
Case 1: $L^{\prime}\nsubseteq L$. Then, $L^{\prime}$ is not a useful reduction of $L$ by item 1 of Definition 6.
Case 2: $L^{\prime}\subseteq L$ and there is an edge $e\in\mathcal{S}(\mathbb{P})^{L}$ such that $L^{\prime}\cap e=\varnothing$. Then, by Definition 15 and statement (8), there is a profile $\delta\in\mathbb{P}$ such that $e=L\cap\mathcal{S}(\delta)$. Thus, $L^{\prime}\cap(L\cap\mathcal{S}(\delta))=\varnothing$ by the case assumption $L^{\prime}\cap e=\varnothing$. Then, $L^{\prime}\cap\mathcal{S}(\delta)=L^{\prime}\cap L\cap\mathcal{S}(\delta)=\varnothing$ by the case assumption $L^{\prime}\subseteq L$. Thus, $L^{\prime}$ is not a useful law in the game $(\mathcal{A},\Delta,\mathbb{P})$ by Lemma 1. Hence, $L^{\prime}$ is not a useful reduction of $L$ by item 1 of Definition 6.
($\Leftarrow$) Suppose $L^{\prime}$ is not a useful reduction of $L$. Then, by item 1 of Definition 6, one of the next cases must be true:
Case 1: $L^{\prime}\nsubseteq L$. Then, $L^{\prime}$ is not a vertex cover of the graph $(L,\mathcal{S}(\mathbb{P})^{L})$ by item 1 of Definition 13.
Case 2: $L^{\prime}$ is not useful. Then, by Lemma 1, there is a profile $\delta\in\mathbb{P}$ such that $L^{\prime}\cap\mathcal{S}(\delta)=\varnothing$. Thus,
$L^{\prime}\cap(L\cap\mathcal{S}(\delta))=L\cap(L^{\prime}\cap\mathcal{S}(\delta))=L\cap\varnothing=\varnothing.$ (10)
Note that $L\cap\mathcal{S}(\delta)\in\mathcal{S}(\mathbb{P})^{L}$ by statement (8) and Definition 15. Hence, $L^{\prime}$ is not a vertex cover of the graph $(L,\mathcal{S}(\mathbb{P})^{L})$ by statement (10) and item 1 of Definition 13. ∎
B.6 Algorithms for the $\mathbf{UL}$ Problems
[…]
2 return $\mathtt{IsVC}$ (graph $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$, set $L$) ;
3
4
5 $\mathtt{IsMiniUL}$(game $(\mathcal{A},\Delta,\mathbb{P})$, set $L$):
6 return $\mathtt{IsMiniVC}$ (graph $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$, set $L$) ;
7
8
9 $\mathtt{AppMinUR}$(game $(\mathcal{A},\Delta,\mathbb{P})$, set $L$):
10 if $\mathtt{IsUL}$($(\mathcal{A},\Delta,\mathbb{P})$, $L$) then
11    return $\mathtt{AppMinVC}$ (graph $(L,\mathcal{S}(\mathbb{P})^{L})$) ;
12
13 return false;
14
Algorithm 1 Algorithms for the $\mathbf{UL}$ Problems
B.7 Approximation Factor of $\mathtt{AppMinUR}$
We first prove the next lemma.
Lemma 8.
For a useful law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$, the tuple $(L,\mathcal{S}(\mathbb{P})^{L})$ is an $|\mathcal{A}|$-graph.
Proof.
Note that $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$ is a $|\mathcal{A}|$-graph by Lemma 7 and $L$ is a vertex cover in the graph $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$ by Theorem 4 and the assumption of the lemma that $L$ is a useful law in the game $(\mathcal{A},\Delta,\mathbb{P})$. Then, the statement of the lemma follows from Definition 15. ∎
Note that, by our assumption, algorithm $\mathtt{AppMinVC}$ in $k$-graphs has an approximation factor of $k$. Meanwhile, by Lemma 8, the subgraph $(L,\mathcal{S}(\mathbb{P})^{L})$ is an $|\mathcal{A}|$-graph. Then, the vertex cover returned in line 7 of Algorithm 1 (say $C$) is at most $|\mathcal{A}|$ times the minimum vertex cover of the input graph $(L,\mathcal{S}(\mathbb{P})^{L})$ (say $\tilde{C}$):
[…]
Meanwhile, by Theorem 5, set $C$ is a useful reduction of $L$ and set $\tilde{C}$ is the minimum useful reduction of $L$. Thus, statement (11) can be interpreted as: the size of the returned set in line 7 of Algorithm 1 is at most $|\mathcal{A}|$ times the size of the minimum useful reduction. Hence, algorithm $\mathtt{AppMinUR}$ in Algorithm 1 is an $|\mathcal{A}|$-approximation of the problem $\mathbf{MinUR}$.
Appendix C Supplementary to Section 4
C.1 Proof of Theorem 6
To support the proof of Theorem 6, we first show the next lemma, which directly follows from Definition 16.
Lemma 9.
For any profile $\bar{\delta}\in\bar{\mathbb{P}}$ in a game $(\bar{\mathcal{A}},\bar{\Delta},\bar{\mathbb{P}})$,
  1. 1.
    if $\bar{\delta}\in\bar{\mathbb{P}}_{1}$, then $\mathcal{S}(\bar{\delta})=\mathcal{S}(\delta)\cup\{p\}$ for a profile $\delta\in\mathbb{P}$;
  2. 2.
    if $\bar{\delta}\in\bar{\mathbb{P}}_{2}$, then $\mathcal{S}(\bar{\delta})=\{d,n\}$ for an action $d\in\bigcup\Delta$;
  3. 3.
    if $\bar{\delta}\in\bar{\mathbb{P}}_{3}$, then $\mathcal{S}(\bar{\delta})=\{n\}$.
Theorem 6. A set $L\subseteq\bigcup\Delta$ is a useful law in a game $(\mathcal{A},\Delta,\mathbb{P})$ if and only if $L$ is a gap-free law in the game $(\bar{\mathcal{A}},\bar{\Delta},\bar{\mathbb{P}})$.
Proof.
Note that $L\subseteq\bigcup\Delta\subsetneq\bigcup\bar{\Delta}$ by Definition 16 and the assumption $L\subseteq\bigcup\Delta$ of the theorem. Then, by Definition 2,
[…]
by Definition 16,
$p,n\notin L,$ (13)
and, by item 1 of Definition 3 and item 2 of Definition 16,
$\bar{\Delta}_{a}^{L}=\begin{cases}\Delta_{a}^{L}\cup\{n\},&\!\text{if }a\in\mathcal{A};\\ \{p,n\},&\!\text{if }a=\gamma.\end{cases}$ (14)
($\Rightarrow$) Suppose law $L$ is useful in the game $(\mathcal{A},\Delta,\mathbb{P})$. Then,
$L\cap\mathcal{S}(\delta)\neq\varnothing\text{ for each profile }\delta\in\mathbb{P}$ (15)
by Lemma 1. Consider an arbitrary profile $\bar{\delta}\in\bar{\mathbb{P}}$ such that
$\bar{\delta}_{\gamma}=p.$ (16)
Then, $\bar{\delta}\in\bar{\mathbb{P}}_{1}$ by item 3 of Definition 16. Thus, $\mathcal{S}(\bar{\delta})=\mathcal{S}(\delta)\cup\{p\}$ for a profile $\delta\in\mathbb{P}$ by item 1 of Lemma 9. Then, $\mathcal{S}(\bar{\delta})\cap L\neq\varnothing$ by statements (15). Hence, $p$ is a safe action of agent $\gamma$ in the law-imposed game $(\bar{\mathcal{A}},\bar{\Delta}^{L},\bar{\mathbb{P}}^{L})$ by statements (14), (16), Lemma 2, and the arbitrariness of $\bar{\delta}$. Therefore, law $L$ is gap-free in the game $(\bar{\mathcal{A}},\bar{\Delta},\bar{\mathbb{P}})$ by Lemma 3 and statement (12).
($\Leftarrow$) Suppose $L$ is not useful in the game $(\mathcal{A},\Delta,\mathbb{P})$. Then, there is a profile $\delta\in\mathbb{P}$ such that $L\cap\mathcal{S}(\delta)=\varnothing$ by Lemma 1. Thus, there is a profile $\bar{\delta}\in\bar{\mathbb{P}}_{1}$ such that $\bar{\delta}_{\gamma}=p$ and $L\cap\mathcal{S}(\bar{\delta})=\varnothing$ by item 3 of Definition 16, item 1 of Lemma 9, and statement (13). Hence, by Lemma 2,
$p$ is not a safe action of agent $\gamma$ (17)
in the law-imposed game $(\bar{\mathcal{A}},\bar{\Delta}^{L},\bar{\mathbb{P}}^{L})$.
Consider an arbitrary agent $a\in\mathcal{A}$ and an arbitrary action $d\in\Delta_{a}^{L}$. Then, $d\notin L$ by item 1 of Definition 3 and there is a profile $\bar{\delta}^{\prime}\in\bar{\mathbb{P}}_{2}$ such that $\bar{\delta}^{\prime}_{a}=d$ and $\mathcal{S}(\bar{\delta}^{\prime})=\{d,n\}$ by item 3 of Definition 16 and item 2 of Lemma 9. Thus, $\mathcal{S}(\bar{\delta}^{\prime})\cap L=\varnothing$ by statement (13). Hence, by Lemma 2, for any agent $a\in\mathcal{A}$ and any action $d\in\Delta_{a}^{L}$,
$d$ is not a safe action of agent $a$ (18)
in the law-imposed game $(\bar{\mathcal{A}},\bar{\Delta}^{L},\bar{\mathbb{P}}^{L})$.
Moreover, there is a profile $\bar{\delta}^{\prime\prime}\in\bar{\mathbb{P}}_{3}$ such that $\bar{\delta}^{\prime\prime}_{a}=n$ for each agent $a\in\bar{\mathcal{A}}$ and $\mathcal{S}(\bar{\delta}^{\prime\prime})=\{n\}$ by item 3 of Definition 16 and item 3 of Lemma 9. Thus, $\mathcal{S}(\bar{\delta}^{\prime\prime})\cap L=\varnothing$ by statement (13). Hence, by Lemma 1 and that $\bar{\delta}^{\prime\prime}\in\bar{\mathbb{P}}_{3}\subseteq\bar{\mathbb{P}}$,
$L$ is not a useful law in the game $(\bar{\mathcal{A}},\bar{\Delta},\bar{\mathbb{P}})$ (19)
and, by Lemma 2, for any agent $a\in\bar{\mathcal{A}}$,
$n$ is not a safe action of agent $a$ (20)
in the law-imposed game $(\bar{\mathcal{A}},\bar{\Delta}^{L},\bar{\mathbb{P}}^{L})$.
In conclusion, by statements (14), (17), (18), and (20), for every agent $a\in\bar{\mathcal{A}}$, every action $d\in\bar{\Delta}_{a}^{L}$ is not a safe action of agent $a$ in the law-imposed game $(\bar{\mathcal{A}},\bar{\Delta}^{L},\bar{\mathbb{P}}^{L})$. Therefore, law $L$ is not gap-free in the game $(\bar{\mathcal{A}},\bar{\Delta},\bar{\mathbb{P}})$ by Lemma 3 and statement (19). ∎
C.2 Proof of Theorem 7
Theorem 7. $\mathbf{MinGFR}$ in a game $(\mathcal{A},\Delta,\mathbb{P})$ is NP-hard to approximate within factor $|\mathcal{A}|-1-\varepsilon$ for any $\varepsilon\!>\!0$ when $|\mathcal{A}|\!\geq\!3$.
Proof.
Suppose the opposite. Then, there is an $\varepsilon>0$ and an algorithm $\mathtt{Alg}$ that approximates $\mathbf{MinGFR}$ with factor $|\mathcal{A}|-1-\varepsilon$ for any game $(\mathcal{A},\Delta,\mathbb{P})$ where $|\mathcal{A}|\geq 3$.
Consider an arbitrary useful law $L$ in an arbitrary game $(\mathcal{A},\Delta,\mathbb{P})$ where $|\mathcal{A}|\geq 2$. We consider the next two steps:
  1. 1.
    compute the game $(\bar{\mathcal{A}},\bar{\Delta},\bar{\mathbb{P}})$ corresponding to the game $(\mathcal{A},\Delta,\mathbb{P})$ by Definition 16;
  2. 2.
    use the algorithm $\mathtt{Alg}$ to get an approximated gap-free reduction $L^{\prime}$ of law $L$ in the game $(\bar{\mathcal{A}},\bar{\Delta},\bar{\mathbb{P}})$.
Note that, by the assumption $|\mathcal{A}|\geq 2$ and Definition 1 of Definition 16,
$|\bar{\mathcal{A}}|=|\mathcal{A}|+1\geq 3.$ (21)
By step 2 above, the assumption that $L$ is a useful law in the game $(\mathcal{A},\Delta,\mathbb{P})$, and Corollary 2,
$L^{\prime}$ is a useful reduction of $L$ in the game $(\mathcal{A},\Delta,\mathbb{P})$. (22)
On the other hand, step 2 above implies that $|L^{\prime}|$ is at most $(|\bar{\mathcal{A}}|-1-\varepsilon)$ times the size of the minimum gap-free reduction of law $L$ in the game $(\bar{\mathcal{A}},\bar{\Delta},\bar{\mathbb{P}})$ by the assumption about the approximation factor of $\mathtt{Alg}$ and statement (21). Meanwhile, the minimum gap-free reduction of law $L$ in the game $(\bar{\mathcal{A}},\bar{\Delta},\bar{\mathbb{P}})$ is the minimum useful reduction of law $L$ in the game $(\mathcal{A},\Delta,\mathbb{P})$ by Corollary 2. Then, by statement (21), the former two facts imply that $|L^{\prime}|$ is at most $(|\mathcal{A}|-\varepsilon)$ times the size of the minimum useful reduction of $L$ in the game $(\mathcal{A},\Delta,\mathbb{P})$. Hence, by statement (22) and the arbitrariness of the useful law $L$ and the game $(\mathcal{A},\Delta,\mathbb{P})$, the above two steps form an algorithm that approximates $\mathbf{MinUR}$ with factor $|\mathcal{A}|-\varepsilon$, which contradicts Theorem 3. ∎
C.3 Proof of Lemma 4
Lemma 4. For a game $(\mathcal{A},\Delta,\mathbb{P})$, an action $d\in\bigcup\Delta$ is safable if and only if $\mathcal{S}(\delta)\neq\{d\}$ for each profile $\delta\in\mathbb{P}$.
Proof.
($\Rightarrow$) Suppose there is a profile $\delta$ such that
$\delta\in\mathbb{P}\text{ and }\mathcal{S}(\delta)=\{d\}.$ (23)
Then, by statement (1),
$\delta_{a}=d\text{ for each agent }a\in\mathcal{A}.$ (24)
Consider an arbitrary agent $a\in\mathcal{A}$ and an arbitrary law $L\subseteq\bigcup\Delta$. Then, one of the next cases must be true:
Case 1: $d\in L$. Then, $d\notin\Delta_{a}^{L}$ by item 1 of Definition 3. Thus, $d$ is not a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$ by Lemma 2.
Case 2: $d\notin L$. Then, $L\cap\mathcal{S}(\delta)=\varnothing$ by the right part of statement (23). Thus, $d$ is not a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$ by the left part of statement (23), statement (24), and Lemma 2.
In conclusion, action $d$ is not safable by Definition 17 and the arbitrariness of $a$ and $L$.
($\Leftarrow$) Suppose $\mathcal{S}(\delta)\neq\{d\}$ for each profile $\delta\in\mathbb{P}$. Let $L$ be the law such that
$L=\bigcup_{\delta\in\mathbb{P},\delta_{a}=d}(\mathcal{S}(\delta)\setminus\{d\}).$ (25)
Then,
$d\notin L.$ (26)
Note that, by the assumption $d\in\bigcup\Delta$ of the lemma and item 2 of Definition 1, there is an agent $a\in\mathcal{A}$ such that $d\in\Delta_{a}$. Thus, by statement (26) and item 1 of Definition 3,
$d\in\Delta_{a}^{L}.$ (27)
If there is no profile $\delta\in\mathbb{P}$ such that $\delta_{a}=d$, then it is vacuously true that $d$ is a safe action of agent $a$ in the game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$ by statement (27) and Lemma 2.
Otherwise, consider an arbitrary profile $\delta^{\prime}\in\mathbb{P}$ such that $\delta^{\prime}_{a}=d$. Then,
$(\mathcal{S}(\delta^{\prime})\setminus\{d\})\cap L=\mathcal{S}(\delta^{\prime})\setminus\{d\}$ (28)
by statement (25), and $\mathcal{S}(\delta^{\prime})\neq\{d\}$ by the ($\Leftarrow$) part assumption. Thus, $\mathcal{S}(\delta^{\prime})\cap L\supseteq(\mathcal{S}(\delta^{\prime})\setminus\{d\})\cap L\neq\varnothing$ by statement (28). Hence, $d$ is a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$ by statement (27), Lemma 2, and the arbitrariness of $\delta^{\prime}$.
In conclusion, action $d$ is safable by Definition 17. ∎
C.4 Proof of Lemma 5
Lemma 5. For a law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$ and an agent $a\in\mathcal{A}$, an action $d\in\bigcup\Delta$ is a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$ if and only if $\in\Delta_{a}^{L}, dd$ is safable in the game $(\mathcal{A},\Delta,\mathbb{P})$, and $L$ is a vertex cover in the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$.
Proof.
($\Rightarrow$) Suppose $d$ is a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$. Then, $d$ is safable in the game $(\mathcal{A},\Delta,\mathbb{P})$ by Definition 17 and $d\in\Delta_{a}^{L}$ by Definition 7. Thus, by item 1 of Definition 3,
$d\in\Delta_{a}\text{ and }d\notin L.$ (29)
Toward proving that $L$ is a vertex cover in the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$, consider an arbitrary profile $\delta\in\mathbb{P}$ such that $\delta_{a}=d$. Then, $L\cap\mathcal{S}(\delta)\neq\varnothing$ by Lemma 2 and the ($\Rightarrow$) part assumption. Thus, $L\cap(\mathcal{S}(\delta)\setminus\{d\})\neq\varnothing$ by the right part of statement (29). Hence, by item 2 of Definition 18 and the arbitrariness of $\delta$,
$L\cap e\neq\varnothing\text{ for each edge $e\in\mathcal{E}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$.}$ (30)
Meanwhile, $L\subseteq\bigcup\Delta$ by Definition 2 and the assumption that $L$ is a law of the game $(\mathcal{A},\Delta,\mathbb{P})$ of the lemma. Then, $L\subseteq\mathcal{V}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$ by the right part of statement (29) and item 1 of Definition 18. Hence, $L$ is a vertex cover of the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$ by Definition 18, statement (30), and item 1 of Definition 13.
$(\Leftarrow)$ Suppose $d\in\Delta_{a}^{L}$ and $L$ is a vertex cover in the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$. Then, $L\cap(\mathcal{S}(\delta)\setminus\{d\})\neq\varnothing$ for each profile $\delta\in\mathbb{P}$ such that $\delta_{a}=d$ by item 2 of Definition 18 and item 1 of Definition 13. Thus, $L\cap\mathcal{S}(\delta)\neq\varnothing$ for each profile $d\in\mathbb{P}$ such that $\delta_{a}=d$. Hence, $d$ is a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$ by Lemma 2 and the $(\Leftarrow)$ part assumption $d\in\Delta_{a}^{L}$. ∎
C.5 Proof of Theorem 9
The next lemma will be used in the later proofs. It follows directly from item 1 of Definition 13.
Lemma 10.
For a vertex cover $C$ in a graph $(V,E)$, every set $C^{\prime}$ such that $C\subseteq C^{\prime}\subseteq V$ is a vertex cover of the graph.
To resolve the complexity in the proof of Theorem 9, we establish Lemma 11 below. It shows that, to verify whether a gap-free law $L$ is minimal, it suffices to check, for each action $d\in L$, whether removing $d$ from law $L$ preserves gap-freeness.
Lemma 11.
A gap-free law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$ is minimal if and only if $L\setminus\{d\}$ is not gap-free for each action $d\in L$.
Proof.
($\Rightarrow$) This part of the statement follows from Definition 10 because $L\setminus\{d\}\subsetneq L$ for each action $d\in L$.
($\Leftarrow$) Suppose a gap-free law $L$ is not minimal. Then,
there is a gap-free law $L^{\prime}\subsetneq L$ (31)
by Definition 10. Thus,
there is an action $d_{0}\in L$ such that $L^{\prime}\subseteq L\setminus\{d_{0}\}$ (32)
and by Lemma 3, one of the next cases must be true:
Case 1: $L^{\prime}$ is useful. Then, $L^{\prime}\cap\mathcal{S}(\delta)\neq\varnothing$ for each profile $\delta\in\mathbb{P}$ by Lemma 1. Thus, $(L\setminus\{d_{0}\})\cap\mathcal{S}(\delta)\neq\varnothing$ for each profile $\delta\in\mathbb{P}$ by statement (32). Then, $L\setminus\{d_{0}\}$ is a useful law in the game $(\mathcal{A},\Delta,\mathbb{P})$ by Lemma 1. Hence, $L\setminus\{d_{0}\}$ is a gap-free law in the game $(\mathcal{A},\Delta,\mathbb{P})$ by Lemma 3, where $d_{0}\in L$ by statement (32).
Case 2: there is an agent $a\in\mathcal{A}$ and a safe action $d_{1}$ of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L^{\prime}},\mathbb{P}^{L^{\prime}})$. Then,
$d_{1}\in\Delta_{a}^{L^{\prime}}\text{ and }L^{\prime}\cap\mathcal{S}(\delta)\neq\varnothing$ (33)
for each profile $\delta\in\mathbb{P}$ such that $\delta_{a}=d_{1}$ by Lemma 2. Thus,
$d_{1}\in\Delta_{a}\text{ and }d_{1}\notin L^{\prime}$ (34)
by item 1 of Definition 3.
Subcase 2.1: $d_{1}\in L$. Then,
$d_{1}\in\Delta_{a}\setminus(L\setminus\{d_{1}\})=\Delta_{a}^{L\setminus\{d_{1}\}}$ (35)
and $L^{\prime}\subseteq L\setminus\{d_{1}\}$ by statements (34), (31), and item 1 of Definition 3. Thus, $(L\setminus\{d_{1}\})\cap\mathcal{S}(\delta)\neq\varnothing$ for each profile $\delta\in\mathbb{P}$ such that $\delta_{a}=d_{1}$ by the right part of statement (33). Then, $d_{1}$ is a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L\setminus\{d_{1}\}},\mathbb{P}^{L\setminus\{d_{1}\}})$ by Lemma 2 and statement (35). Hence, law $L\setminus\{d_{1}\}$ is a gap-free by Lemma 3, where $d_{1}\in L$ by the subcase assumption.
Subcase 2.2: $d_{1}\notin L$. Then, $d_{1}\notin L\setminus\{d_{0}\}$. Thus, by the left part of statement (34) and item 1 of Definition 3,
$d_{1}\in\Delta_{a}\setminus(L\setminus\{d_{0}\})=\Delta_{a}^{L\setminus\{d_{0}\}}.$ (36)
Note that $(L\setminus\{d_{0}\})\cap\mathcal{S}(\delta)\neq\varnothing$ for each profile $\delta\in\mathbb{P}$ such that $\delta_{a}=d_{1}$ by statements (33) and (32). Then, $d_{1}$ is a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L\setminus\{d_{0}\}},\mathbb{P}^{L\setminus\{d_{0}\}})$ by Lemma 2 and statement (36). Hence, law $L\setminus\{d_{0}\}$ is a gap-free by Lemma 3, where $d_{0}\in L$ by statement (32). ∎
Theorem 9. A gap-free law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$ is minimal if and only if all of the following statements are true:
  1. 1.
    if $L$ is a vertex cover of the graph $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$, then $L$ is a minimal vertex cover in this graph;
  2. 2.
    for each agent $a\in\mathcal{A}$ and each safable action $d\in\Delta_{a}^{L}$, if $L$ is a vertex cover in graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$, then $L$ is a minimal vertex cover in this graph;
  3. 3.
    for each agent $a\in\mathcal{A}$ and each safable action $d\!\in\!\Delta_{a}\!\cap\!L$, the set $L\!\setminus\!\{d\}$ is not a vertex cover in graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$.
Proof.
($\Rightarrow$) If item 1 is false, then, by item 2 of Definition 13, there is a set $L^{\prime}$ such that
$L^{\prime}\subsetneq L$ (37)
being a vertex cover of the graph $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$. Thus, $L^{\prime}$ is a gap-free law in the game $(\mathcal{A},\Delta,\mathbb{P})$ by Theorem 8. Hence, $L$ is not minimal-gap-free by Definition 10 and statement (37).
If item 2 is false, then, by item 2 of Definition 13, there is an agent $a$ and a safable action $d\in\Delta_{a}^{L}$, and a vertex cover $L^{\prime}$ in the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$ such that
$L^{\prime}\subsetneq L.$ (38)
Thus, $d\in\Delta_{a}^{L}=\Delta_{a}\setminus L\subseteq\Delta_{a}\setminus L^{\prime}=\Delta_{a}^{L^{\prime}}$ by item 1 of Definition 3. Then, action $d$ is a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L^{\prime}},\mathbb{P}^{L^{\prime}})$ by Lemma 5. Hence, law $L^{\prime}$ is gap-free in the game $(\mathcal{A},\Delta,\mathbb{P})$ by Lemma 3. Then, $L$ is not minimal-gap-free by statement (38) and Definition 10.
If item 3 is false, then there is an agent $a$ and a safable action $d\in\Delta_{a}\cap L$ such that
$L\setminus\{d\}$ is a vertex cover in the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$. (39)
Thus, $d\in\Delta_{a}$ and $d\in L$. Then,
$L\setminus\{d\}\subsetneq L$ (40)
and $d\in\Delta_{a}\setminus(L\setminus\{d\})=\Delta_{a}^{L\setminus\{d\}}$ by item 1 of Definition 3. Hence, action $d$ is a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L\setminus\{d\}},\mathbb{P}^{L\setminus\{d\}})$ by Lemma 5, statement (39), and that $d$ is safable. Then, law $L\setminus\{d\}$ is gap-free in the game $(\mathcal{A},\Delta,\mathbb{P})$ by Lemma 3. Therefore, $L$ is not minimal-gap-free by Definition 10 and statement (40).
($\Leftarrow$) Suppose gap-free law $L$ is not minimal. Then, by Lemma 11, there is an action $d$ such that
$d\in L$ (41)
and $L\setminus\{d\}$ is gap-free. Thus,
$L\setminus\{d\}\subsetneq L\subseteq\bigcup\Delta$ (42)
and, by Theorem 8, one of the next cases must be true:
Case 1: $L\setminus\{d\}$ is a vertex cover of the graph $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$. Then, $L$ is a vertex cover of the graph $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$ by Lemma 10 and statement (42). Meanwhile, $L$ is not a minimal vertex cover of the graph $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$ by the case assumption, statement (42), and item 2 of Definition 13. Therefore, item 1 of the theorem is false.
Case 2: there is an agent $a\in\mathcal{A}$ and a safable action $d^{\prime}\in\Delta_{a}^{L\setminus\{d\}}$ such that
$L\setminus\{d\}$ is a vertex cover of the graph $\mathcal{H}^{a,d^{\prime}}_{(\mathcal{A},\Delta,\mathbb{P})}$. (43)
Note that $\Delta_{a}^{L}\subseteq\Delta_{a}^{L\setminus\{d\}}$ by by item 1 of Definition 3. Then, by the case assumption $d^{\prime}\in\Delta_{a}^{L\setminus\{d\}}$, one of the next subcases must be true:
Subcase 2.1: $d^{\prime}\in\Delta_{a}^{L}$. Then, $d^{\prime}\notin L$ by item 1 of Definition 3. Thus, $L\subseteq\bigcup\Delta\setminus\{d^{\prime}\}=\mathcal{V}^{a,d^{\prime}}_{(\mathcal{A},\Delta,\mathbb{P})}$ by statement (42) and item 1 of Definition 18. Hence, $L$ is a vertex cover of the graph $\mathcal{H}^{a,d^{\prime}}_{(\mathcal{A},\Delta,\mathbb{P})}$ by statement (43) and Lemma 10. Meanwhile, $L$ is not a minimal vertex cover of the graph $\mathcal{H}^{a,d^{\prime}}_{(\mathcal{A},\Delta,\mathbb{P})}$ by statements (43), (42), and item 2 of Definition 13. Therefore, item 2 of the theorem is false by the subcase assumption $d^{\prime}\in\Delta_{a}^{L}$ and the case assumption $d^{\prime}$ is safable.
Subcase 2.2: $d^{\prime}\in\Delta_{a}^{L\setminus\{d\}}\setminus\Delta_{a}^{L}$. It is worth noting that $\Delta_{a}^{L\setminus\{d\}}\setminus\Delta_{a}^{L}=\Delta_{a}\cap\{d\}$ by item 1 of Definition 3. Then, $d^{\prime}=d\in\Delta_{a}\cap L$ by the subcase assumption and statement (41). Thus, item 3 of the theorem is false by statement (43) and the case assumption $d^{\prime}$ is safable. ∎
C.6 Proof of Theorem 10
To improve the readability, we first show several lemmas that will support the proof of Theorem 10.
Lemma 12.
For a graph $(V,E)$ and two sets $C^{\prime},C\subseteq V$, the next two statements are equivalent:
  1. 1.
    $C^{\prime}\subseteq C$ and set $C^{\prime}$ is a vertex cover of the graph $(V,E)$.
  2. 2.
    $C$ is a vertex cover of the graph $(V,E)$ and $C^{\prime}$ is a vertex cover of the subgraph $(C,E^{C})$.
Proof.
($1\Rightarrow 2$) Note that set $C$ is a vertex cover of the graph $(V,E)$ by the assumption $C\subseteq V$ of the lemma, item 1 above, and Lemma 10.
Moreover, $C^{\prime}\cap e\neq\varnothing$ for each edge $e\in E$ by item 1 of Definition 13 and item 1 above. Then, $C^{\prime}\cap(C\cap e)=C^{\prime}\cap C\cap e=C^{\prime}\cap e\neq\varnothing$ for each egde $e\in E$ by the assumption $C^{\prime}\subseteq C$ in item 1 above. Hence, set $C^{\prime}$ is a vertex cover of the graph $(C,E^{C})$ by Definition 15, item 1 of Definition 13, and the assumption $C^{\prime}\subseteq C$ in item 1 above.
($1\Leftarrow 2$) Note that, by item 2 above, item 1 of Definition 13, and Definition 15,
$C^{\prime}\subseteq C\subseteq V$ (44)
and $C^{\prime}\cap(C\cap e)\neq\varnothing$ for each egde $e\in E$. Thus, $C^{\prime}\cap e\supseteq C^{\prime}\cap C\cap e\neq\varnothing$ for each egde $e\in E$. Hence, $C^{\prime}$ is a vertex cover of the graph $(V,E)$ by item 1 and Definition 13 and statement (44). ∎
Lemma 13.
For any two laws $L^{\prime},L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$, the next two statements are equivalent:
  1. 1.
    $L^{\prime}\subseteq L$ and law $L^{\prime}$ is useful in the game $(\mathcal{A},\Delta,\mathbb{P})$.
  2. 2.
    $L$ is a vertex cover in the graph $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$ and $L^{\prime}$ is a vertex cover in the subgraph $(L,\mathcal{S}(\mathbb{P})^{L})$.
Proof.
The statement of the lemma follows from Definition 2, Theorem 4, and Lemma 12. ∎
Lemma 14.
For any two laws $L^{\prime},L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$, an agent $a\in\mathcal{A}$, and a safable action $d\in\Delta_{a}^{L}$, the next two statements are equivalent:
  1. 1.
    $L^{\prime}\subseteq L$ and $d$ is a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L^{\prime}},\mathbb{P}^{L^{\prime}})$.
  2. 2.
    $L$ is a vertex cover in the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$ and $L^{\prime}$ is a vertex cover in the subgraph $(L,(\mathcal{E}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})})^{L})$.
Proof.
Note that, by the assumption $L^{\prime},L$ are two laws of the game $(\mathcal{A},\Delta,\mathbb{P})$ and Definition 2,
$L^{\prime},L\subseteq\bigcup\Delta.$ (45)
($1\Rightarrow 2$) By Lemma 5 and item 1 above, $L^{\prime}$ is a vertex cover in the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$. Then, item 2 above is true by Lemma 12, statement (45), and the assumption $L^{\prime}\subseteq L$ in item 1 above.
($1\Leftarrow 2$) By Lemma 12 and item 2 above, $L^{\prime}\subseteq L$ and
$L^{\prime}$ is a vertex cover in the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$. (46)
Then, $d\in\Delta_{a}^{L}\subseteq\Delta_{a}^{L^{\prime}}$ by item 1 of Definition 3 and the assumption $d\in\Delta_{a}^{L}$ of the lemma. Thus, $d$ is a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L^{\prime}},\mathbb{P}^{L^{\prime}})$ by Lemma 5, statement (46), and the assumption of the lemma that action $d$ is safable. ∎
Lemma 15.
For any two laws $L^{\prime},L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$, an agent $a\in\mathcal{A}$, and a safable action $d\in\Delta_{a}\cap L$, the next two statements are equivalent:
  1. 1.
    $L^{\prime}\subseteq L$ and $d$ is a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L^{\prime}},\mathbb{P}^{L^{\prime}})$.
  2. 2.
    $L\setminus\{d\}$ is a vertex cover in the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$ and $L^{\prime}$ is a vertex cover in the subgraph $(L\setminus\{d\},(\mathcal{E}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})})^{L\setminus\{d\}})$.
Proof.
Note that, by the assumption that $L^{\prime},L$ are two laws of the game $(\mathcal{A},\Delta,\mathbb{P})$ and Definition 2,
$L^{\prime},L\subseteq\bigcup\Delta.$ (47)
($1\Rightarrow 2$) Note that $d\in\Delta_{a}^{L^{\prime}}$ by the second part of item 1 above and Lemma 2. Then, $d\notin L^{\prime}$ by item 1 of Definition 3. Thus, $d\in L\setminus L^{\prime}$ by the assumption $d\in\Delta_{a}\cap L$ of the lemma. Hence, by the first part of item 1 above,
$L^{\prime}\subseteq L\setminus\{d\}.$ (48)
Moreover, $L\setminus\{d\}\subseteq(\bigcup\Delta)\setminus\{d\}$ by statement (47). Then, by statement (48) and item 1 of Definition 18,
$L^{\prime}\subseteq L\setminus\{d\}\subseteq\mathcal{V}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}.$ (49)
Besides, by Lemma 5 and the second part of item 1 above,
[…]
Then, item 2 above follows from statements (49) and (50) by Lemma 12 and Definition 18.
($1\Leftarrow 2$) By item 1 of Definition 13 and item 2 above,
$L^{\prime}\subseteq L\setminus\{d\}\subseteq\mathcal{V}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}.$ (51)
Then, $L^{\prime}\subseteq L$ and $d\notin L^{\prime}$. Thus, by item 1 of Definition 3 and the assumption $d\in\Delta_{a}\cap L$ of the lemma,
$d\in\Delta_{a}\setminus L^{\prime}=\Delta_{a}^{L^{\prime}}.$ (52)
Moreover, by Lemma 12 and statement (51), item 2 above implies that $L^{\prime}$ is a vertex cover in the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$. This, together with statement (52) and the assumption of the lemma that action $d$ is safable, further implies that $d$ is a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L^{\prime}},\mathbb{P}^{L^{\prime}})$ by Lemma 5. ∎
Theorem 10. For a gap-free law $L$ in a game $(\mathcal{A},\Delta,\mathbb{P})$, a law $L^{\prime}$ is a gap-free reduction of $L$ if and only if at least one of the following statements is true:
  1. 1.
    $L$ is a vertex cover of the graph $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$ and $L^{\prime}$ is a vertex cover in the subgraph $(L,\mathcal{S}(\mathbb{P})^{L})$;
  2. 2.
    there is an agent $a\in\mathcal{A}$ and a safable action $d\in\Delta_{a}^{L}$ such that $L$ is a vertex cover in the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$ and $L^{\prime}$ is a vertex cover in the subgraph $(L,(\mathcal{E}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})})^{L})$;
  3. 3.
    there is an $a\!\in\!\mathcal{A}$ and a safable action $d\!\in\!\Delta_{a}\!\cap\!L$ such that $L\!\setminus\!\{d\}$ is a vertex cover in graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$ and $L^{\prime}$ is a vertex cover in the subgraph $(L\!\setminus\!\{d\},(\mathcal{E}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})})^{L\setminus\{d\}})$.
Proof.
($\Rightarrow$) Suppose $L^{\prime}$ is a gap-free reduction of $L$. Then,
$L^{\prime}\subseteq L\subseteq\bigcup\Delta$ (53)
and $L^{\prime}$ is gap-free in the game $(\mathcal{A},\Delta,\mathbb{P})$ by Definition 2 and item 1 of Definition 11. Thus, by Lemma 3, one of the next cases must be true:
Case 1: $L^{\prime}$ is useful in the game $(\mathcal{A},\Delta,\mathbb{P})$. Then, item 1 of the theorem is true by statement (53) and Lemma 13.
Case 2: there is an agent $a\in\mathcal{A}$ and a safe action $d$ of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L^{\prime}},\mathbb{P}^{L^{\prime}})$. Then,
$d\in\Delta_{a}^{L^{\prime}}$ (54)
by Lemma 2, and
action $d$ is safable (55)
by Definition 17. Note that $\Delta_{a}^{L^{\prime}}\supseteq\Delta_{a}^{L}$ by statement (53) and item 1 of Definition 3. Then, by statement (54), one of the next two subcases must be true:
Subcase 2.1: $d\in\Delta_{a}^{L}$. Then, by Lemma 14, item 2 of the theorem follows from statements (53), (55), and the Case 2 assumption.
Subcase 2.2: $d\in\Delta_{a}^{L^{\prime}}\setminus\Delta_{a}^{L}$. Note that
$\Delta_{a}^{L^{\prime}}\setminus\Delta_{a}^{L}=\Delta_{a}\cap(L\setminus L^{\prime})\subseteq\Delta_{a}\cap L.$
Then, $d\in\Delta_{a}\cap L$ by the subcase assumption. Thus, by Lemma 15, item 3 of the theorem follows from statement (53), (55), and the Case 2 assumption.
($\Leftarrow$) Note that, by item 1 of Definition 11, it suffices to show that, if any of the three items of the theorem is true, then $L^{\prime}\subseteq L$ and $L^{\prime}$ is a gap-free law in the game $(\mathcal{A},\Delta,\mathbb{P})$.
If item 1 is true, then $L^{\prime}\subseteq L$ and $L^{\prime}$ is useful in the game $(\mathcal{A},\Delta,\mathbb{P})$ by Lemma 13 and the assumption that $L,L^{\prime}$ are both laws in the game. Thus, $L^{\prime}$ is a gap-free law in the game $(\mathcal{A},\Delta,\mathbb{P})$ by Lemma 3.
If item 2 is true, then $L^{\prime}\subseteq L$ and $d$ is a safe action of agent $a$ in the game $(\mathcal{A},\Delta^{L^{\prime}},\mathbb{P}^{L^{\prime}})$ by Lemma 14 and the assumption that $L,L^{\prime}$ are both laws in the game. Thus, $L^{\prime}$ is a gap-free law in the game $(\mathcal{A},\Delta,\mathbb{P})$ by Lemma 3.
If item 3 is true, then $L^{\prime}\subseteq L$ and $d$ is a safe action of agent $a$ in the game $(\mathcal{A},\Delta^{L^{\prime}},\mathbb{P}^{L^{\prime}})$ by Lemma 15 and the assumption that $L,L^{\prime}$ are both laws in the game. Thus, $L^{\prime}$ is a gap-free law in the game $(\mathcal{A},\Delta,\mathbb{P})$ by Lemma 3. ∎
C.7 Approximation Factor of $\mathtt{AppMinGFR}$
First, observe that all graphs in the statement of Theorem 10 are $|\mathcal{A}|$-graphs by Lemma 7, Definition 15, and Definition 18. This means, all the $\mathtt{AppMinVC}$ input graphs in lines 35, 40, and 43 of Algorithm 2 are $|\mathcal{A}|$-graphs.
Moreover, by Theorem 10, the minimum gap-free reduction (say $R$) must be a minimum vertex cover in at least one of the $\mathtt{AppMinVC}$ input graphs (say the minimum vertex cover $C$ of graph $G$) in lines 35, 40, and 43 of Algorithm 2:
$|R|=|C|.$ (56)
Then, by the assumption about the approximation factor of $\mathtt{AppMinVC}$, the execution of $\mathtt{AppMinVC}$ on graph $G$ returns a vertex cover (say $C^{\prime}$) whose size is at most $|\mathcal{A}|$ times the size of the minimum vertex cover $C$ in graph $G$:
$|C^{\prime}|\leq|\mathcal{A}|\times|C|.$ (57)
Moreover, by lines 33, 35, 40, and 43 of Algorithm 2, the size of the final output in line 44 of Algorithm 2 is no bigger than $|C^{\prime}|$:
$|\mathit{output}|\leq|C^{\prime}|.$ (58)
Hence, by statements (56), (57), and (58),
$|output|\leq|\mathcal{A}|\times|R|,$
which means the size of the final output of $\mathtt{AppMinGFR}$ in Algorithm 2 is at most $|\mathcal{A}|$ times the size of the minimum gap-free reduction. Therefore, $\mathtt{AppMinGFR}$ in Algorithm 2 is an $|\mathcal{A}|$-approximation of the problem $\mathbf{MinGFR}$.
C.8 Algrithms for the $\mathbf{GFL}$ Problems
(See the next page.)
[…]
2 if $d\notin\bigcup\Delta$ then
3    return false;
4
5 for each profile $\delta\in\mathbb{P}$ do
6    if $\mathcal{S}(\delta)=\{d\}$ then
7       return false;
8
9
10 return true;
11
12
13 $\mathtt{IsGFL}$(game $(\mathcal{A},\Delta,\mathbb{P})$, set $L$):
14 if $\mathtt{IsVC}$ ($(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$, $L$) then //item 1 of Theorem 8
15    return true;
16
17 for each agent $a\in\mathcal{A}$ and each action $d\in\Delta_{a}$ do //item 2 of Theorem 8
18    if $d\notin L$ and $\mathtt{IsSafable}$ ($(\mathcal{A},\Delta,\mathbb{P})$,$d$) then
19       if $\mathtt{IsVC}$ (graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$, set $L$) then
20          return true;
21
22
23
24 return false ;
25
26
27 $\mathtt{IsMiniGFL}$(game $(\mathcal{A},\Delta,\mathbb{P})$, set $L$):
28 if not $\mathtt{IsGFL}$ ($(\mathcal{A},\Delta,\mathbb{P})$, $L$) then //Check if set $L$ is a gap-free law
29    return false;
30
31 if $\mathtt{IsVC}$ ($(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$, $L$) and not $\mathtt{IsMiniVC}$ ($(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$, $L$) then //item 1 of Theorem 9
32    return false;
33
34 for each agent $a\in\mathcal{A}$ and each action $d\in\Delta_{a}$ do
35    if $\mathtt{IsSafable}$ ($(\mathcal{A},\Delta,\mathbb{P})$,$d$) then
36       if $d\notin L$ then //item 2 of Theorem 9
37          if $\mathtt{IsVC}$ ($\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$, $L$) and not $\mathtt{IsMiniVC}$ ($\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$, $L$) then
38             return false;
39
40
41       else //item 3 of Theorem 9
42          if $\mathtt{IsVC}$ ($\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$, $L\setminus\{d\}$) then
43             return false;
44
45
46
47
48 return true ;
49
50
51 $\mathtt{AppMinGFR}$(game $(\mathcal{A},\Delta,\mathbb{P})$, set $L$):
52 if not $\mathtt{IsGFL}$ ($(\mathcal{A},\Delta,\mathbb{P})$, $L$) then //Check if set $L$ is a gap-free law
53    return false;
54
55 $\mathit{output}\leftarrow L$;
56 if $\mathtt{IsVC}$ ($(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$, $L$) then //item 1 of Theorem 10
57    $\mathit{output}\leftarrow\min\{\mathit{output},\mathtt{AppMinVC}(L,\mathcal{S}(\mathbb{P})^{L})\}$;
58
59 for each agent $a\in\mathcal{A}$ and each action $d\in\Delta_{a}$ do
60    if $\mathtt{IsSafable}$ ($(\mathcal{A},\Delta,\mathbb{P})$,$d$) then
61       if $d\notin L$ then //item 2 of Theorem 10
62          if $\mathtt{IsVC}$ ($\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$, $L$) then
63             $\mathit{output}\leftarrow\min\{\mathit{output},\mathtt{AppMinVC}(L,(\mathcal{E}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})})^{L})\}$;
64
65
66       else //item 3 of Theorem 10
67          if $\mathtt{IsVC}$ ($\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$, $L\setminus\{d\}$) then
68             $\mathit{output}\leftarrow\min\{\mathit{output},\mathtt{AppMinVC}(L\setminus\{d\},(\mathcal{E}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})})^{L\setminus\{d\}})\}$;
69
70
71
72
73 return $\mathit{output}$ ;
74
Algorithm 2 Algorithms for the $\mathbf{GFL}$ Problems